### v1

# Symplectic Hodge Theory on Lie Algebroids

Abstract.
We explore the natural analogues of the Brylinksi condition, Strong Lefschetz condition, and $d\delta$-lemma in Symplectic Geometry originally explored by Brylinksi, Mathieu, Yan, and Guillemin to the Symplectic Lie Algebroid case. The equivalence of the three conditions is re-established as a purely algebraic statement along with a primitive notion of the $d\delta$-lemma shown establsihed by Tseng, Yau, and Ho. We then show that the natural analogues of these in the Lie Algebroid setting holds as well with examples given
1. Introduction
1.1. Background
Hodge theory for Symplectic Manifolds was introduced by Ehresmann and Liberman and rediscovered by Brylinski [10.4310/jdg/1214442161]; Brylinksi conjectured that on a Symplectic Manifold $(M,\omega)$ of dimension $2m$ every deRham cohomology class admits a symplectic harmonic representative. Mathieu [Mathieu] and Yan [YAN1996143] independently showed that the Brylinksi conjecture holds true if and only if $(M^{2m},\omega)$ satisfies the Strong Lefschetz Property, that is for all $0\leq k\leq m$, the map
$[L]^{k}:H^{m-k}(M)$ $\rightarrow H^{m+k}(M)$
$[\alpha]$ $\mapsto[\omega^{k}\wedge\alpha]$
is surjective. This result was then improved by Merkulov in [merkulov1998formality], and Guillemin in [Guilleminddlemma] by showing that these two conditions are equivalent to the symplectic $d\delta$-lemma in the case that $M$ was a compact manifold using Poincaré duality. In a key departure from traditional Riemannian Hodge theory the Symplectic Laplacian is no longer an elliptic operator as it vanishes identically, and as such the results that heavily makeup the theory no longer work. Instead we primarily work with the observation that in the presence of a symplectic form, the space of differential forms on a smooth manifold admits the structure of an $\mathfrak{sl}_{2}(\mathbb{R})$-module. This “Symplectic Hodge Theory” has been extended to new settings in work by Lin [equivariantddlemma] and others, which require reestablishing many conditions in specific geometric settings. In this paper we work towards streamlining this by proving analogous statements about modules in a relevant module category, and applying them to Symplectic Lie Algebroids, a minimal smooth setting for the notions involved to make sense. We also generalize the operators $\partial_{+},\partial_{-}$ of Tseng and Yau [tseng2012cohomologyhodgetheorysymplectic] to this strictly algebraic setting (which we rename $\operatorname{\mathfrak{d}}$ and $\operatorname{\overline{\mathfrak{d}}}$ respectively), though we do not explore the related Laplacians, or Appelli and Bott-Chern cohomologies corresponding to certain combinations of these operators. We then relate the original trio of equivalences back to $\partial_{+}\partial_{-}$-lemma of [HoSym] by reestablishing the equivalence of this to the original triple in our algebraic framework rather than the original topological framework considered. Through this generalization we are also able to lift the compactness condition, though it was free to assume in the classical case of the tangent bundle by Poincaré Duality and standard Algebraic Topological facts relating to the relevant conditions.
1.2. Results
We define a category of modules $\mathscr{C}_{n}$ over a suitable Lie Superalgebra with properties that axiomatize the algebra of differential forms on a Symplectic Manifold much in the spirit of [MR1480723].
Definition 1.
Let $\mathfrak{g}$ be the 5 dimensional Lie Superalgebra with basis $e,f,h,d,\delta$ with $e,f,h$ degrees $2,-2,0$ respectively and $d,\delta$ in degrees $1,-1$ respectively all subject to the following relations
$[e,f]$ $=h$ $[e,h]$ $=-2e$ $[f,h]$ $=2f$
$[e,d]$ $=0$ $[f,d]$ $=\delta$ $[h,d]$ $=d$
$[e,\delta]$ $=d$ $[f,\delta]$ $=0$ $[h,\delta]$ $=-\delta$
$[d,d]$ $=0$ $[d,\delta]$ $=0$ $[\delta,\delta]$ $=0$
where all commutators are graded commutators. As a Lie Superalgebra this is isomorphic to the $5$-dimensional Orthosymplectic Lie Superalgebra
It is well known that the space of forms of any symplectic manifold admits the structure of a $\mathfrak{g}$-module where the even part of the algebra corresponds to the standard $\mathfrak{sl}_{2}$ structure, $d$ corresponds to the deRham differential, and $\delta$ is defined in [10.4310/jdg/1214442161]. Since the space of forms admits finitely many degree, 00 through $2n$, we want to restrict our attention to modules for which $h$ acts with finitely many eigenvalues. Moreover, since $h$ acts on the form integrally and via a diagonalizable action, we want to restrict to modules for which this occurs as well.
Definition 2.
For $n\geq 0$, let $\mathscr{C}_{n}$ denote the full subcategory of $\mathfrak{g}{\operatorname{\mathbf{-mod}}}$ such that $h$ acts integrally via diagonalizable action with finitely many eigenvalues between $-n$ and $n$
Objects of this category are also referred to as “Modules of Finite $h$-type” or “Modules of Finite $h$-spectrum” in the literature. In the case that the underlying geometric structure is Kähler, or more generally has the Hard Lefschetz Property, the module enjoys similar properties and is called a “Lefschetz Module”.
Theorem.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. Then the following are equivalent
  1. (1)
    The map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $i$
  2. (2)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ is a quasi-isomorphism
  3. (3)
    The $d\delta$-lemma holds for $V$
  4. (4)
    The $\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma holds for $V$
After introducing the notion of the $d\delta$-lemma holding weakly in section 2.4, we also weaken this to the following theorem
Theorem.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. Then the following are equivalent
  1. (1)
    The $d\delta$-lemma holds for $V$ up to degree $s$
  2. (2)
    The map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $|i|\geq s$
  3. (3)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ is a quasi-isomorphism in degree $|i|\geq s$
Applying the first of these two theorems to a Symplectic Lie Algebroid with representation we have the following theorem:
Theorem.
Let $(\mathcal{A}\rightarrow M,\rho,\omega)$ be a Symplectic Lie Algebroid of fiber dimension $2m$, with representation $(E,\nabla)$. Then the following are equivalent:
  1. (1)
    The map $[L]^{k}:H^{m-k}_{\mathcal{A}}(M,E)\rightarrow H^{m+k}_{\mathcal{A}}(M,E)$ given by $[\omega]^{k}\wedge-$ is an isomorphism for all $k$
  2. (2)
    The inclusion of complexes $(\ker({}^{\mathcal{A}}\nabla^{*}),{}^{\mathcal{A}}\nabla)\hookrightarrow(\Omega_{\mathcal{A}}^{\bullet}(M,E),{}^{\mathcal{A}}\nabla)$ is a quasi-isomorphism
  3. (3)
    The $(\mathcal{A},E)$-$d\delta$-lemma holds for $\mathcal{A}$
  4. (4)
    The $(\mathcal{A},E)$-$\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma holds for $\mathcal{A}$
Choosing $E$ to be the trivial line bundle on $M$, $\mathcal{A}=TM$, and with the standard deRham differential $d$ as the connection we recover the main results of [YAN1996143], [Mathieu], and [HoSym]. In section $2$ we introduce the relevant Lie Superalgebra and associated module category and go on to establish the four results equivalence as a purely algebraic fact as well as reestablish the notion of the “weak” Lefschetz property as seen in [fernández2005ddeltalemmaweaklylefschetzsymplectic]. In section $3$ we introduce Symplectic Lie Algebroids, and define operators on the space of $\mathcal{A}$-forms analogous to those on a Symplectic Manifold. Section $4$ contains examples of the theorems proven. First, the classical example $KT^{4}$ The Kodaira-Thurston manifold, fits well into the framework as it should. Next, using Ovando’s [ovando2004dimensional] classification of symplectomorphism classes of $4$-dimensional Lie Algebras, we discuss which admit the structure of a Lefschetz Module. Moving on we consider six-dimensional Nilmanifolds, and discuss which of the 26 isomorphism classes admitting symplectic forms admit Lefschetz maps that are isomorphisms. We then discuss $E$-manifolds as studied by Miranda and Scott [Miranda_2020], and finally Kähler Lie Algebroids.
2. Lefschetz Modules
2.1. The Lie Superalgebra
Proposition 2.1.
If $V\in\mathrm{Ob}(\mathscr{C}_{n})$, then $V$ admits a decomposition
$V=\bigoplus_{i=-n}^{n}V_{i}$
where $h$ acts on $V_{i}$ as scaling by $i$, i.e. $i\cdot\text{id}_{V_{i}}$
Lemma 2.1.
Every cyclic $V\in\mathrm{Ob}(\mathscr{C}_{n})$ is finite dimensional
Proof.
Let $v$ generate $V$ as a $U(\mathfrak{\mathfrak{sl}_{2}(\mathbb{R})})$-module, and decompose $v$ into it’s components $v=\sum v_{i}$ for $v_{i}\in V_{i}$. It suffices to show that the submodule each $v_{i}$ generate is finite dimensional. By the Poincaré-Birkhoff-Witt theorem, this module is spanned by the vectors
$\langle v_{i}\rangle=\operatorname{span}_{m,r,k}\{e^{r}f^{m}h^{k}v_{i}\}$
Since $h$ acts by scaling on $v_{i}$, it suffices to check then on
$\langle v_{i}\rangle=\operatorname{span}_{m,r}\{e^{r}f^{m}v_{i}\}$
Now, since there are only finitely many eigenvalues, we have that $r,m\leq n$. From this we can note that there are finitely many combinations of $m$ and $r$ such that
$-n\leq-2m+2r+i\leq n$
Thus there are finitely many generators for $\langle v_{i}\rangle$
Now, we’d like to be able to talk about primitive elements as normal in $\mathfrak{sl}_{2}$ representation theory.
Definition 3.
Suppose $V\in\mathrm{Ob}(\mathscr{C}_{n})$ with decomposition $V=\bigoplus_{i}V_{i}$. We call a vector $v_{i}\in V_{i}$ primitive if $fv_{i}=0$ or equivalently $e^{i+1}v_{i}=0$
Using this definition, we reprove a common structure theorem for $\mathfrak{sl}_{2}$-modules
Theorem 2.2.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. Then every $v\in V$ can be written as a finite sum
$v=\sum e^{i}v_{i}$
where the $v_{i}$ are primitive
Proof.
Start by replacing $V$ with $\langle v\rangle$. This is again an object of $\mathscr{C}_{n}$, and as it’s cyclic must be finite dimensional. We can then decompose it into a finite sum of irreducibles, and write $v$ as it’s components in these, so it suffices to show elements in these irreducibles is of the form desired. Suppose $v\in V_{-n}$ and $u\in V_{-n+1}$, both of which are primitive by degree considerations. Now, since $fe^{k}v=\lambda_{1}e^{k-1}v$ and $fe^{k}u=\lambda_{2}e^{k-1}u$ for nonzero constants $\lambda_{i}$, we know that they span a nonzero submodule together. Since we’re in the irreducible case, this must be the entire module. Moreover, they all correspond to distinct eigenvalues of $h$, so they must be linearly independent, i.e. they form a basis of $V$. ∎
Another useful fact is the following classical result
Proposition 2.2.
Suppose $V\in\mathrm{Ob}(\mathscr{C}_{n})$ with decomposition $V=\bigoplus_{i}V_{i}$. Then the map $e^{i}:V_{-i}\rightarrow V_{i}$ is a bijection
Proof.
This follows from classical $\mathfrak{sl}_{2}(\mathbb{R})$ representation theory ∎
As well as the following bracket relations
Lemma 2.3.
$[e^{i},\delta]=ie^{i-1}d$
Proof.
We proceed by induction, the base case is given as definition. Suppose this holds for the $i-1$ case, then we have that
$e^{i}\delta-\delta e^{i}$ $=e(\delta e^{i-1}+(i-1)e^{i-2}d)-e^{i}\delta$
$=e\delta e^{i-1}+(i-1)e^{i-1}d-\delta e^{i}$
$=(\delta e+d)e^{i-1}+(i-1)e^{i-1}d-\delta e^{i}$
$=ie^{i-1}d$
Lemma 2.4.
$[h,e^{i}]=2ie^{i}$
Proof.
This holds from definition for $i=1$, so suppose it holds for $i-1$. Then we have that
$he^{i}-e^{i}h$ $=(e^{i-1}h+2(i-1)e^{i-1})e-e^{i}h$
$=e^{i-1}he+2(i-1)e^{i}-e^{i}h$
$=e^{i-1}(eh+2e)+2(i-1)e^{i}-e^{i}h$
$=2ie^{i}$
We also want to bring the notion of the $\star$ involution present in classical Symplectic Hodge Theory:
Definition 4.
On any $V\in\mathrm{Ob}(\mathscr{C}_{n})$, define the “star operator” $\star:V_{i}\rightarrow V_{-i}$ to be the involution such that
$\delta|_{V_{i}}=(-1)^{i+1}\star d\star$
Now, in the spirit of classical deRham cohomology we arrive at the following definitions:
Definition 5.
Let $V\in\mathscr{C}_{n}$, and $v\in V$. Then we define the following
  1. (1)
    $v$ is “closed” if $v\in\ker(d)$
  2. (2)
    $v$ is “exact” if $v\in\operatorname{im}(d)$
  3. (3)
    $v$ is “coclosed” if $v\in\ker(\delta)$
  4. (4)
    $v$ is “coexact” if $v\in\operatorname{im}(\delta)$
More generally we can adopt the language of harmonicity as follows
Definition 6.
An element $v\in V$, where $V\in\mathscr{C}_{n}$, we say that $v$ is “harmonic” if $v\in\ker(d)\cap\ker(\delta)$. The submodule of harmonic elements of $V$ will be denoted $\widehat{V}$
Proposition 2.3.
For $V\in\mathrm{Ob}(\mathscr{C}_{n})$, $\widehat{V}$ is a $\mathfrak{g}$-submodule of $V$.
Proof.
This is clearly a vector subspace and closed under the action of $d$ and $\delta$. If $v\in\widehat{V}$, then $dev=edv=0$ and $\delta ev=e\delta v-dv=0$. Likewise $dfv=fdv+\delta v=0$ and $\delta fv=f\delta v=0$. The closedness under $h$ follows from $e$ and $f$
One can regard $V$ as a bi-differential $\mathbb{Z}$-graded complex after writing $V$ out as
[…]
Definition 7.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$ with decomposition $V=\bigoplus_{i}V_{i}$. The cohomology of $V$ regarded as the above complex in $d$ at the $i$-th position is denoted $H^{i}(V,d)$
Definition 8.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$ with decomposition $V=\bigoplus_{i}V_{i}$. The homology of $V$ regarded as the above complex in in $\delta$ at the $i$-th position is denoted $H_{i}(V,\delta)$
Lemma 2.5.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$, and $v\in V$. If $v$ is closed, then $\star v$ is coclosed. Similarly, if $v$ is coclosed, then $\star v$ is is closed.
Proof.
Suppose that $dv=0$. Then we have that
$\delta(\star v)=\pm\star d\star\star v=\pm\star dv=0$
Now, suppose $\delta v=0$. Then we have that
$\delta v$ $=0$
±$\pm\star d\star v$ $=0$
±$\pm d\star v$ =0=0=$\star$ 0=0= ⋆ 0 = 0
$\implies d(\star v)$ $=0$
Lemma 2.6.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$, and $v\in V$. If $v$ is exact, then $\star v$ is coexact. If $v$ is coexact, then $\star v$ is exact
Proof.
Suppose that $v$ is exact, that is $v=dw$. Then we have that
$\star v=\star dw=\star d\star\star w=\pm\delta(\star w)$
Similarly, suppose $v=\delta w$. Then we have
$\star v=\star\delta w=\pm\star\star d\star w=d(\pm\star w)$
Theorem 2.7.
For $V\in\mathrm{Ob}(\mathscr{C}_{n})$, there is an isomorphism
$H^{i}(\ker(\delta),d)\cong H^{-i}(\ker(d),\delta)$
Given by the $\star$ map
Proof.
Define a map
$\varphi:\widehat{V}_{i}/d(V_{i-1}\cap\ker(\delta))\rightarrow\widehat{V}_{-i}/\delta(V_{i+1}\cap\ker(d))$
Given by sending $[v]\mapsto[\star v]$. This is well-defined by the preceding lemmas, and is surjective as for any class $[v]$, it has preimage $[\star v]$. To see injectivity, suppose that $\varphi[v]$ vanishes, then by definition we have
$\star v=\delta w=\pm\star d\star w$
For some $w\in\ker(d)\cap V_{i+1}$, and the sign depending on the grading. Applying $\star$ to the above relation we have
$v=\pm d\star w=d(\pm\star w)$
Which means that $[v]$ was the zero class. The only possible point of confusion is why ±$\pm\star w$ is in the kernel of $\delta$, but this again follows from the preceding lemmas ∎
2.2. Equivalence of Surjectivity
The following theorem was established independently by Yan and Mathieu.
Theorem 2.8 (Yan [YAN1996143], Mathieu [Mathieu]).
Let $(M^{2n},\omega)$ be a symplectic manifold. Then the following are equivalent:
  1. (1)
    The strong Lefschetz map $[\omega^{k}]:H^{n-k}_{dR}(M)\rightarrow H^{n+k}_{dR}(M)$ is surjective
  2. (2)
    Every deRham Cohomology class has has a (symplectic) harmonic representative
We can rephrase this purely as an algebraic equivalence: Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$, then the following are equivalent:
  1. (1)
    For all $0\leq i\leq n$, the map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is surjection
  2. (2)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ induces a surjection in cohomology $H^{i}(\ker(\delta),d)\rightarrow H^{i}(V,d)$
Lemma 2.9.
$[e^{i}]$ carries harmonic harmonic representatives to harmonic representatives
Proof.
Suppose $[v_{-i}]\in H^{-i}(V,d)$ has a harmonic representative $[v_{-i}^{0}]$. We claim that $[e^{i}v_{-i}^{0}]$ is a harmonic representative of $[e^{i}v_{-i}]$. First, note that they’re cohomologous as if $v_{-i}=v_{-i}^{0}+d\tau_{-i-1}$, then we have that
$e^{i}v_{-i}=e^{i}v_{-i}^{0}+e^{i}d\tau_{-i-1}=e^{i}v_{-i}^{0}+de^{i}\tau_{-i-1}$
Moreover, we have that $e^{i}v_{-i}^{0}$ is harmonic since
$\delta e^{i}v_{-i}=e^{i}\delta v_{-i}-ie^{i-1}dv_{-i}=0$
The following is essentially a rephrasing of Yan’s original proof [YAN1996143]:
Theorem 2.10 (D. Yan).
Suppose $V\in\mathrm{Ob}(\mathscr{C}_{n})$, then the following statements are equivalent
  1. (1)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ induces a surjection in cohomology $H^{i}(\ker(\delta),d)\rightarrow H^{i}(V,d)$
  2. (2)
    For all $0\leq i\leq n$, the map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is a surjection
Proof.
Suppose condition $1$ holds, that is the maps $H^{i}(\ker(\delta),d)\rightarrow H^{i}(V,d)$ are surjections, or equivalently we can say we have surjections $\widehat{V}_{i}\rightarrow H^{i}(V,d)$. We also know by classical $\mathfrak{sl}_{2}$-representation theory that $e^{i}:V_{-i}\rightarrow V_{i}$ is a bijection, and this restricts to a bijection on $\widehat{V}_{-i}\rightarrow\widehat{V}_{i}$ since the harmonic elements make up a $\mathfrak{g}$-submodule of $V$. Together these give us the diagram
[…]
Where the top and vertical arrows are surjections, implying the bottom map is surjective. This is the exact map in condition 2. Now, suppose that $[e^{i}]$ is surjective for all $i$. We’d like to show that the inclusion map $(\ker(\delta),d)\hookrightarrow(V,d)$ is a surjection in cohomology, i.e. given any cohomology class $[v]\in H^{\bullet}(V,d)$, we can replace it by a class $[w]$ where $w\in\ker(\delta)$. First, we claim that it suffices to check this on cohomologies in nonpositive degree. To see this, fix $[v_{i}]\in H^{i}(V,d)$ for $i>0$. Then as $[e^{i}]$ as surjective, we have that $[v_{i}]=[e^{i}][w_{-i}]$ for some $[w_{-i}]\in H^{-i}(V,d)$, and we know by Propostion 2.9 that $[e^{i}]$ will carry a harmonic representative of $[w_{-i}]$ to a harmonic representative of $[v_{i}]$. Now, we claim that for $i>0$, $H^{-i}(V,d)$ decomposes
$H^{-i}(V,d)=\operatorname{im}([e])+\ker([e^{i+1}])$
Fix a class $[v_{-i}]\in H^{-i}(V,d)$ and consider $[e^{i+1}v_{-i}]\in H^{i+2}(V,d)$. Since the $[e^{i}]$ are surjective, we have that there is class $[\beta_{-i-2}]\in H^{-i-2}(V,d)$ such that $[e^{i+2}\beta_{-i-2}]=[e^{i+1}v_{-i}]$. Now, we can (trivially) write
$[v_{-i}]=([v_{-i}-e\beta_{-i-2}])+[e]([\beta_{-i-2}])$
It’s clear that the second summand is in the image of $[e]$, we need only to check that the first is in the kernel of $[e^{i+1}]$. This follows from the definition as
$[e^{i+1}]([v_{-i}-e\beta_{-i-2}])=[e^{i+1}v_{-i}-e^{i+2}\beta_{-i-2}]=[e^{i+1}v_{-i}-e^{i+1}v_{-i}]=[0]$
We now proceed to prove the claim by induction. First, note that every class in $H^{-n}(V,d)$ and $H^{-n+1}(V,d)$ is harmonic. If $[v_{-n}]\in H^{-n}(V,d)$, then $\delta v_{-n}\equiv 0$ trivially, and so $v_{-n}$ is harmonic. If $[v_{-n+1}]\in H^{-n+1}(V,d)$, then we have that
$\delta v_{-n+1}=[d,f]v_{-n+1}=dfv_{-n+1}-fdv_{-n+1}=0$
Since $fv_{-n+1}\in V_{-n-1}\equiv 0$, and $v_{-n+1}$ is closed. So we only need to check on $-n+2\leq i\leq 0$, and the $-n+1$ degree serves as the base case. Now, suppose the statement holds for $H^{-i-1}(V,d)$ and we want to show that it holds on $H^{-i}(V,d)$. Fix $[w_{-i}]\in H^{-i}(V,d)$. By the above decomposition we have that
$[\beta_{-i}]=[\alpha_{-i}]+[e]([\beta_{-i-2}])$
For $[\beta_{-i-2}]\in H^{-i-2}(V,d)$ and $[\alpha_{-i}]\in H^{-i}(V,d)\cap\ker([e^{i+1}])$. By inductive hypothesis $[\beta_{-i-2}]$ possesses a harmonic representative, and $[e]$ carries harmonic representatives into harmonic representatives, so we can freely assume $[e]([\beta_{-i-2}]$ is harmonic. We only need to show that $[\alpha_{-i}]$ admits a harmonic representative. By definition of $[\alpha_{-i}]$ we have that $d\alpha_{-i}=0$ and $e^{i+1}\alpha_{-i}=d\beta_{i+1}$ for some $\beta_{i+1}\in V_{i+1}$. Now, by the bijectivity of $e^{i+1}:V_{-i-1}\rightarrow V_{i+1}$, there exists $\gamma_{-i-1}\in V_{-i-1}$ such that $\beta_{i+1}=e^{i+1}\gamma_{-i-1}$. We claim that $\alpha_{-i}-d\gamma_{-i-1}$ is a harmonic representative of $[\alpha_{-i}]$. It’s clear that $[\alpha_{-i}]=[\alpha_{-i}-d\gamma_{-i-1}]$, and that $\alpha_{-i}-d\gamma_{-i-1}\in\ker(d)$, we only need to show that it’s in $\ker(\delta)$. Since $\delta=[d,f]$, this reduces to showing that $\alpha_{-i}-d\gamma_{-i-1}\in\ker(f)$, i.e. is primitive. Primitivity here is equivalent to showing that $\alpha_{-i}-d\gamma_{-i-1}\in\ker(e^{i+1})$. Computing we have that
$e^{i+1}(\alpha_{-i}-d\gamma_{-i-1})=e^{i+1}\alpha_{-i}-d(e^{i+1}\gamma_{-i-1})=e^{i+1}\alpha_{-i}-d\beta_{i+1}=e^{i+1}\alpha_{-i}-e^{i+1}\alpha_{-i}=0$
2.3. Equivalence of Triples
In the previous subsection we’ve shown that the following two statements are equivalent for $V\in\mathrm{Ob}(\mathscr{C}_{n})$:
  1. (1)
    The map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is a surjection
  2. (2)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ induces a surjection in cohomology
Which in the original paper(s), Mathieu claims is equivalent to a third condition: The $d\delta$-lemma.
Definition 9.
The $d\delta$-lemma holds for $V\in\mathrm{Ob}(\mathscr{C}_{n})$ if the following equalities hold
$\operatorname{im}(d)\cap\ker(\delta)=\operatorname{im}(\delta)\cap\ker(d)=\operatorname{im}(d\delta)$
This is equivalent to the above two if we replace surjection with isomorphism in both claims; in the original setting of the tangent bundle of a compact manifold these are equivalent by Poincare duality. The first of the equivalences - that $d\delta$-lemma is equivalent to the inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ follows from the following:
Theorem 2.11.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. Then the $d\delta$-lemma is equivalent to the inclusion $(\ker(\delta),d)\rightarrow(V,d)$ inducing an isomorphism in cohomology
Proof.
See lemma 5.4.1 in [manin1998constructions]
Proposition 2.4.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$, and suppose that the $d\delta$-lemma is satisfied. Then $[e^{k}]:H^{-k}(V,d)\rightarrow H^{k}(V,d)$ is an isomorphism for all $k$
Proof.
Since $d\delta$-lemma implies that there are harmonic representatives, we have that that $[e^{k}]$ is surjective by the prior work. This also allows us to check the validity of this isomorphism by restricting to $[e^{k}]:H^{-k}(\ker(\delta),d)\rightarrow H^{k}(\ker(\delta),d)$. So, suppose that $[\alpha_{-k}]\in\ker([e^{k}])$. Then we have that $e^{k}\alpha_{-k}=d\tau_{k-1}$ for some $\tau_{k-1}\in V_{k-1}\cap\ker(\delta)$. Then, note that
$\delta(e^{k}(\alpha_{-k}))=\delta d\tau_{k-1}=-d\delta\tau_{k-1}=0$
So, the form $e^{k}\alpha_{-k}\in\operatorname{im}(d)\cap\ker(\delta)$, so by the $d\delta$-lemma we have that there is a $\rho_{k}$ such that
$e^{k}\alpha_{-k}=d\delta\rho_{k}$
Now, since $e^{k}$ is a bijection for all $k$, let $\phi^{k}$ denote it’s inverse. Then we have that
$\alpha_{-k}=\phi^{k}(d\delta\rho_{k})=-\phi^{k}\delta d\rho_{k}$
Now, since $[e^{k},\delta]=ke^{k-1}d$, we have that
$\phi^{k}\delta=\delta\phi^{k}-k\phi^{k}e^{k-1}d\phi^{k}$
Meaning that
$a_{-k}$ $=-(\delta\phi^{k}-k\phi^{k}e^{k-1}d\phi^{k})(d\rho_{k})$
$=-\delta\phi^{k}d\rho_{k}+k\phi^{k}e^{k-1}\phi^{k}d^{2}\rho_{k}$
$=-\delta\phi^{k}d\rho_{k}$
$=d\delta\phi^{k}\rho_{k}$
So $\alpha_{-k}\in\operatorname{im}(d)$ for $\delta\phi^{k}\rho_{k}\in\ker(\delta)$
Proposition 2.5.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. If $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $i$, then $\ker(d)\cap\operatorname{im}(\delta)=\operatorname{im}(d)\cap\operatorname{im}(\delta)$.
Proof.
What we want to show is that for a general module element $v\in V$, we have that $d\delta v=0$ implies that $\delta v=d\phi$ for some $\phi$. First, we claim that it suffices to check this on primitive elements. Towards this, suppose that the statement holds on primitive elements, and fix arbitrary $v\in V$. We can decompose $v$ into primitive elements, i.e. $v=\sum e^{i}v_{i}$ where the $v_{i}$ are primitive. Now, since $d\delta v=0$, we have that
$0=d\delta v=\sum d\delta e^{i}v_{i}=\sum d(e^{i}\delta-ie^{-i1}d)v_{i}=\sum e^{i}d\delta v_{i}$
Since this sum is direct, this forces $e^{i}d\delta v_{i}=0$ for all $i$. Since $e_{i}$ is acting on the negative weights in this sum, it acts injectively, so $d\delta v_{i}=0$ for all $i$. By hypothesis, we have then that there exist $\phi_{i}$ such that $\delta v_{i}=d\phi_{i}$, and thus
$\delta v$ $=\sum\delta e^{i}v_{i}$
$=\sum e^{i}\delta v_{i}-ie^{i-1}dv_{i}$
$=\sum e^{i}d\phi_{i}-ie^{i-1}dv_{i}$
$=d\left(\sum e^{i}\phi_{i}-ie^{i-1}v_{i}\right)$
Now, to see that this holds on primitive form we proceed as follows. Suppose $v_{-i}\in V_{-i}$ is primitive and $d\delta v_{-i}=0$. Then we have that
$e^{i+1}\delta v_{-i}=\delta e^{i+1}v_{-i}-(i+1)e^{i}dv_{-i}$
However, $e^{i+1}v_{-i}=0$ since $v_{-i}$ is primitive, and we’re left with
$e^{i+1}\delta v_{-i}=d(-(i+1)e^{i}v_{-i})$
Rephrased this says that $[\delta v_{-i}]\in\ker[e^{i+1}]$, which by the Strong Lefschetz Property is an isomorphism, meaning that $[\delta v_{-i}]$ is the zero class in cohomology, so $\delta v_{-i}=d\phi$ for some $\phi\in V_{-i-2}$. The only degree where this argument does not hold is $V_{-n}$, but for any $v\in V_{-n}$, $\delta v\equiv 0$, so it’s trivially in the image of $d$, i.e. $\delta v=0=d(0)$. ∎
Proposition 2.6.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$, and suppose that $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $i$. Then $\ker(d)\cap\operatorname{im}(\delta)=\ker(\delta)\cap\operatorname{im}(d)$
Proof.
From the preceding proposition, we have that $\ker(d)\cap\operatorname{im}(\delta)=\operatorname{im}(d)\cap\operatorname{im}(\delta)\subset\ker(\delta)\cap\operatorname{im}(d)$. So we need to show that $\operatorname{im}(d)\cap\ker(\delta)\subset\operatorname{im}(\delta)\cap\ker(d)$. So suppose $\alpha\in\operatorname{im}(d)\cap\ker(\delta)$, that is $\alpha=d\beta$ and $\delta\alpha=0$. Then we have
$0=\delta\alpha=\pm\star d\star\alpha$
Where sign depends on degree, but is not of importance for this proof. Since $\star$ is an isomorphism, we have that $d\star\alpha=0$,i.e. $\star\alpha\in\ker(d)$. Now using exactness we have
$\star\alpha=\star d\beta=\pm\star d\star\star\beta=\pm\delta\star\beta$
Thus $\star\alpha\in\ker(d)\cap\operatorname{im}(\delta)$. By the preceding proposition yet again we have then that $\star\alpha\in\operatorname{im}(d)\cap\operatorname{im}(\delta)$, and so $\star\alpha=d\rho$. Using this we then have that
$\alpha=\star\star\alpha=\star d\rho=\star d\star\star\rho=\delta(\pm\star\rho)$
Giving us $\alpha\in\operatorname{im}(\delta)$
Proposition 2.7.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$, and suppose that $[e]^{i}:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is surjective for all $i$. Then $\operatorname{im}(d)\cap\operatorname{im}(\delta)=\operatorname{im}(d\delta)$
Proof.
Since the two operators anticommute, the $\supset$ direction is clear, and we need only to prove the forward inclusion which we’ll prove by induction. First note that the statement trivially holds elements in $V_{n}$ or $V_{-n}$ since if $v\in\operatorname{im}(\delta)\cap V_{n}$, then $v=\delta 0=0$, and likewise for $V_{-n}$. Our base case is on $V_{n-1}$, so choose $v_{n-1}\in V_{n-1}$, and let
$v_{n-1}=d\gamma_{n-2}=\delta w_{n}$
Now, since $dw_{n}$ is trivially zero it represents a cohomology class, and so we can choose a harmonic representative of said class giving us $w_{n}=w_{n}^{0}-d\tau_{n-1}$ where $w_{n}^{0}$ is harmonic. We then have that
$v_{n-1}$ $=\delta w_{n}$
$=\delta(w_{n}^{0}-d\tau_{n-1})$
$=-\delta d\tau_{n-1}$
$=d\delta\tau_{n-1}$
Now, suppose this holds for $V_{k}$ for some $-n<k<n$, and fix $v_{k-1}\in V_{k-1}\cap\operatorname{im}(d)\cap\operatorname{im}(\delta)$. By definition we have that
$v_{k-1}=d\gamma_{k-2}=\delta\beta_{k}$
From this, define $\xi_{k+1}=d\beta_{k}$, and observe that
$\delta\xi_{k+1}=\delta d\beta_{k}=-d(\delta\beta_{k})=-d^{2}\gamma_{k-2}=0$
So $\xi_{k+1}\in\operatorname{im}(d)\cap\ker(\delta)$. By the remark preceding this proposition we have that $\operatorname{im}(d)\cap\ker(\delta)=\operatorname{im}(d)\cap\operatorname{im}(\delta)$, which by inductive hypothesis extends to $\operatorname{im}(d)\cap\ker(\delta)=\operatorname{im}(d)\cap\operatorname{im}(\delta)=\operatorname{im}(d\delta)$, so $\xi_{k+1}=d\delta\phi_{k+1}$. Then note that $d(\beta_{k}-\delta\phi_{k+1})=0$, so it admits a harmonic representative, i.e. $\beta_{k}-\delta\phi_{k+1}=\beta_{k}^{0}-d\tau_{k}$ for some $\tau_{k}\in V_{k}$. Then applying the definition of $v_{k-1}$ we have
$v_{k-1}$ $=\delta\beta_{k}$
$=\delta(\beta_{k}^{0}-d\tau_{k}+\delta\phi_{k+1})$
$=-\delta d\tau_{k}$
$=d\delta\tau_{k}$
At this point, we have enough to establish the equivalence of the triple: $d\delta$-lemma, Strong Lefschetz property, and Brylinksi property. We’ve already proven all the of the relevant pieces, but for clarity state them here in one place
Theorem 2.12.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. The the following conditions are equivalent:
  1. (1)
    The $d\delta$-lemma holds for $V$
  2. (2)
    The map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $i$
  3. (3)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ is a quasi-isomorphism
Proof.
We have that the first and third conditions are equivalent by Theorem 2.11. Then we have that the first implies the second by Proposition 2.4, and the reverse direction follows from the Proposition 2.5, 2.6, and 2.7. ∎
In light of this theorem, we make the following definition
Definition 10.
If $V\in\mathrm{Ob}(\mathscr{C}_{n})$ satisfies any of the above equivalent conditions in Theorem 2.12 we call $V$ a “Lefschetz Module”
Corollary 2.12.1.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$ be a Lefschetz module. Then $H^{*}(V,d)$ is an $\mathfrak{sl}_{2}(\mathbb{R})$-module
Proof.
We can see from the $\mathfrak{g}$ relations that
$[e,d]=0\qquad[f,d]=\delta\qquad[h,d]=d$
So on $H^{*}(V,d)$, we see that $e$ and $h$ pass to cohomology, i.e. it’s always a $B$-module. The only possible problem lies in well-definedness of $f$. Suppose $[v]=[w]\in H^{*}(V,d)$, that is $v=w+d\phi$ for some $\phi\in V$. Then we have that
$f[v]=[fv]=[f(w+d\phi)]=f[w]+[fd\phi]=f[w]+[(df+\delta)\phi]=f[w]+[\delta\phi]$
Now $\delta\phi\in\operatorname{im}(\delta)\cap\ker(d)$, so since $V$ is Lefschetz, there exists some $\varphi$ such that $\delta\phi=d\delta\varphi$ so
$f[v]=f[w]+[d\delta\varphi]=f[w]$
2.4. Weakly Lefschetz Modules
There are examples on manifolds where the Lefschetz condition is only partially met, i.e. the maps $[L]^{k}$ are isomorphisms for some $k$, but not all $0\leq k\leq n$
Definition 11.
Given $V\in\mathrm{Ob}(\mathscr{C}_{n})$, we say that $V$ is “$s$-Lefschetz” if $e^{k}$ is a isomorphism for $0\leq s\leq k\leq n$
The same theorems from before hold in this setting as well, but only on the portions on the modules for which the maps $e^{k}$ are isomorphisms.
Theorem 2.13.
Suppose $V\in\mathrm{Ob}(\mathscr{C}_{n})$ , then the following statements are equivalent
  1. (1)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ induces a surjection in cohomology $H^{i}(\ker(\delta),d)\rightarrow H^{i}(V,d)$ in degrees $|i|\geq s$
  2. (2)
    For all $0\leq s\leq i\leq n$, the map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is a surjection
Proof.
The forward direction follows in the same exact way it does in the proof of Theorem 2.10. The reverse direction followed by induction that crucially used the fact that we could decompose $H^{-i}(V,d)$ for $i>0$. This decomposition was the only part of the proof that relied on the reverse direction, and the rest of the proof holds if we only assume that the $[e^{i}]$ are surjective from $V_{-n}$ up to $V_{-s}$. ∎
The upgraded equivalence also holds under the added assumption that $e^{i}$ is in fact an isomorphism
Definition 12.
Given a module $V\in\mathrm{Ob}(\mathscr{C}_{n})$, we say that $V$ has the “$d\delta$-lemma up to degree $s$” if
$\operatorname{im}(d)\cap\ker(\delta)$ $=\operatorname{im}(d\delta)=\operatorname{im}(\delta)\cap\ker(\delta)$ on $V_{i}\text{ for }-n\leq i\leq s$
$\operatorname{im}(d)\cap\ker(\delta)$ $=\operatorname{im}(d\delta)$ on $V_{s+1}$
for $-n\leq s\leq-1$
Note that by $\star$ duality if this equality holds on $V_{i}$ for $-n\leq i\leq-s$ it then also holds in $V_{i}$ for $s\leq i\leq n$. From this observation we’ll only prove statements about it in non-positive degree, as the duality will extend the results to positive degrees.
Theorem 2.14.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. Then the following are equivalent
  1. (1)
    The $d\delta$-lemma holds for $V$ up to degree $s$
  2. (2)
    The map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $|i|\geq s$
  3. (3)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ is a quasi-isomorphism in degree $|i|\geq s$
Proof.
The proof will follow similarly to the original. The first and third condition are equivalent by [manin1998constructions], whose argument works in each degree. The first implies the second using Proposition 2.4 and acknowledging the proof is symmetric in it’s degree arguments about the zero weight. Finally, to see that the second condition implies the first, note that proposition 2.5 and proposition 2.6 are degree agnostic, and proposition 2.7 is done by induction, which we can once again stop early when the hypothesis fail in degree $-s$ and achieve the result ∎
2.5. Primitive Equivalence
Much of this section is rephrasing of work in Chung-I Ho’s thesis [HoSym] and Tseng’s paper [tseng2012cohomologyhodgetheorysymplectic]
Definition 13.
Suppose $V\in\mathrm{Ob}(\mathscr{C}_{n})$ with decomposition $\bigoplus_{k}V_{k}$. Then the space of primitive elements of degree $k$ will be denoted $\mathcal{P}_{k}=\ker(f)\cap V_{k}$
Throughout the rest of this section, fix $V\in\mathrm{Ob}(\mathscr{C}_{n})$, so that we may speak of elements $v_{i}\in\operatorname{\mathcal{P}}_{i}$ freely.
Lemma 2.15.
Suppose that $w_{i}\in\mathcal{P}_{i}$. Then the primitive decomposition of $dw_{i}$ only has two terms
Proof.
First, note that
$f^{2}dw_{i}=f(df+\delta)w_{i}=f\delta w_{i}=\delta fw_{i}=0$
As $w_{i}$ is primitive. Taking a primitive decomposition of $dw_{i}$ we have
$dw_{i}=\sum_{k\geq 0}e^{k}v_{i+1-2k}$
So applying $f^{2}$ we have
$0=\sum_{k\geq 0}f^{2}e^{k}v_{i+1-2k}$
Where $v_{i+1-2k}\in\mathcal{P}_{i+1-2k}$ Which means for all $k\geq 2$, we had to have had that $e^{k}v_{i+1-2k}=0$, but as $e^{i}$ acts injectively we have that $v_{i+1-2k}=0$ for these $k$, leaving us with
$dw_{i}=v_{i+1}+ev_{i-1}$
Definition 14.
Given $v_{i}\in\mathcal{P}_{i}$ with $dv_{i}=v_{i+1}+ev_{i-1}$, define
$\operatorname{\mathfrak{d}}v_{i}$ $=v_{i+1}$
$\operatorname{\overline{\mathfrak{d}}}v_{i}$ $=v_{i-1}$
so on $\operatorname{\mathcal{P}}_{i}$ we have $dv_{i}=v_{i+1}+ev_{i-1}=\operatorname{\mathfrak{d}}v_{i}+e\operatorname{\overline{\mathfrak{d}}}v_{i}$
These are the same operators as $\partial_{+}$ and $\partial_{-}$ as introduced in [tseng2012cohomologyhodgetheorysymplectic]. These extend to the entire module by taking primitive decomposition of any element $v\in V$ , i.e. if $v=\sum v_{i}$ for $v_{i}\in V_{i}$, and $v_{i}=\sum_{k\geq 0}e^{k}v_{i-2k}$ then we have that
$\operatorname{\mathfrak{d}}v$ $=\sum_{i}\sum_{k\geq 0}e^{k}\operatorname{\mathfrak{d}}v_{i-2k}$
$\operatorname{\overline{\mathfrak{d}}}v$ $=\sum_{i}\sum_{k\geq 0}e^{k}\operatorname{\overline{\mathfrak{d}}}v_{i-2k}$
Lemma 2.16.
We have that $[\operatorname{\mathfrak{d}},e]=[\operatorname{\overline{\mathfrak{d}}},e]=[\operatorname{\mathfrak{d}},\operatorname{\mathfrak{d}}]=[\operatorname{\overline{\mathfrak{d}}},\operatorname{\overline{\mathfrak{d}}}]=0$ where all are graded commutators, as well as
$e(\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}})=-e(\operatorname{\overline{\mathfrak{d}}}\operatorname{\mathfrak{d}})$
On $V_{n-1},V_{n}$ and $[\operatorname{\mathfrak{d}},\operatorname{\overline{\mathfrak{d}}}]=0$ on $V_{k}$ for $k<n-1$
Proof.
This follows the same as it did in [tseng2012cohomologyhodgetheorysymplectic], Lemma 2.5 ∎
Proposition 2.8.
For $v_{i}\in\operatorname{\mathcal{P}}_{i}$, we have that $f\operatorname{\mathfrak{d}}v_{i}=0$
Proof.
Since $v_{i}$ is primitive, we have that $[f,\operatorname{\mathfrak{d}}]v_{i}=f\operatorname{\mathfrak{d}}v_{i}$. By the Jacobi identity we have that
$[f,\operatorname{\mathfrak{d}}]=\frac{1}{2}[h,[\operatorname{\mathfrak{d}},f]]+[f,\operatorname{\mathfrak{d}}]$
Thus we have to have that $h$ and $[\operatorname{\mathfrak{d}},f]$ commute. Checking this on $v_{i}$ gives us that
$(i-1)[\operatorname{\mathfrak{d}},f]v_{i}=i[\operatorname{\mathfrak{d}},f]v_{i}$
Forcing $[f,\operatorname{\mathfrak{d}}]v_{i}=0$
Proposition 2.9.
For $v_{i}\in\operatorname{\mathcal{P}}_{i}$ $\delta v_{i}=(1-i)\operatorname{\overline{\mathfrak{d}}}v_{i}$
Proof.
Using the definition of $\delta$ we have
$\delta v_{i}=fdv_{i}-dfv_{i}=fdv_{i}=f\operatorname{\mathfrak{d}}v_{i}+fe\operatorname{\overline{\mathfrak{d}}}v_{i}$
We know the first summand vanishes by the preceding lemma, and on the second we have that
$fe\operatorname{\overline{\mathfrak{d}}}v_{i}=(ef-h)\operatorname{\overline{\mathfrak{d}}}v_{i}=-h\operatorname{\overline{\mathfrak{d}}}v_{i}=-(i-1)\operatorname{\overline{\mathfrak{d}}}v_{i}$
Lemma 2.17.
On $\operatorname{\mathcal{P}}_{i}$, we have that $d\delta=(1-i)\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$
Proof.
Using the preceding lemmas we have
$d\delta v_{i}=(i-1)d\operatorname{\overline{\mathfrak{d}}}v_{i}=(i-1)(\operatorname{\mathfrak{d}}+e\operatorname{\overline{\mathfrak{d}}})\operatorname{\overline{\mathfrak{d}}}v_{i}=(1-i)\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}v_{i}$
Definition 15.
We say that $v\in V$ is “symphonic” if $\operatorname{\mathfrak{d}}v=\operatorname{\overline{\mathfrak{d}}}v=0$
This was previously called “primitive harmonic” in the literature, see [HoSym], however to avoid the phrase “harmonic primitive harmonic” we introduce this terminology.
Lemma 2.18.
$v\in V$ is harmonic if and only if each term of it’s primitive decomposition is symphonic
Proof.
Let $v=v_{i}$ with $v_{i}=\sum_{k\geq 0}e^{k}w_{i-2k}$ for $w_{i-2k}\in\operatorname{\mathcal{P}}_{i-2k}$. Then if each $w_{i-2k}$, since $d=\operatorname{\mathfrak{d}}+e\operatorname{\overline{\mathfrak{d}}}$ it’s clear that $v$ is closed. Using the commutation relation given in Lemma 2.3, we have
$\delta v$ $=\sum_{i}\sum_{k\geq 0}\delta e^{k}w_{i-2k}$
$=\sum_{i}\sum_{k\geq 0}(e^{k}\delta-ke^{k-1}d)w_{i-2k}$
$=\sum_{i}\sum_{k\geq 0}e^{k}(1-i)\operatorname{\overline{\mathfrak{d}}}w_{i-2k}$
$=0$
Now, suppose that $v=\sum_{i}v_{i}$ for $v_{i}\in V_{i}$ is harmonic. this occurs if and only if each $v_{i}$ is harmonic, which occurs if and only if the term in the primitive decomposition of each $v_{i}$ is harmonic, so it suffices to chow this for a primitive harmonic element. Towards this, suppose $v\in\widehat{\operatorname{\mathcal{P}}_{i}}$. Then
$0=\delta v=(i-1)\operatorname{\overline{\mathfrak{d}}}v$
So $v\in\ker(\operatorname{\overline{\mathfrak{d}}})$. Moreover we have
$0=dv=\operatorname{\mathfrak{d}}v+e\operatorname{\overline{\mathfrak{d}}}v=\operatorname{\mathfrak{d}}v$
Thus $v\in\ker(\operatorname{\mathfrak{d}})$
Definition 16.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. We say that $V$ satisfies the $\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma if the following equalities hold
$\operatorname{im}(\operatorname{\mathfrak{d}})\cap\ker(\operatorname{\overline{\mathfrak{d}}})\cap\operatorname{\mathcal{P}}_{k}$ $=\operatorname{im}(\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}})\cap\operatorname{\mathcal{P}}_{k}\qquad-n+1\leq k\leq-1$
$\operatorname{im}(\operatorname{\mathfrak{d}})\cap\ker(\operatorname{\overline{\mathfrak{d}}})\cap\operatorname{\mathcal{P}}_{-n}$ $=\operatorname{im}(\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}})\cap\operatorname{\mathcal{P}}_{-n}$
$\operatorname{im}(\operatorname{\overline{\mathfrak{d}}})\cap\ker(\operatorname{\mathfrak{d}})\cap\operatorname{\mathcal{P}}_{0}$ $=\operatorname{im}(\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}})\cap\operatorname{\mathcal{P}}_{0}$
Lemma 2.19.
The $d\delta$-lemma holds for $V\in\mathrm{Ob}(\mathscr{C}_{n})$ if the $\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma does
Proof.
It suffices to check this on each weight, so fix $v_{i}\in\widehat{V_{i}}$ and first suppose that $v_{i}\in\operatorname{im}d$ so that $v_{i}=dw_{i-1}$. Then taking primitive decompositions of both we have
$v_{i}=\sum_{k\geq 0}e^{k}v_{i-2k}\qquad w_{i-1}=\sum_{k\geq 0}e^{k}w_{i-1-2k}$
Combining this with the fact that $v_{i}=dw_{i-1}$ we have
$\sum_{k\geq 0}e^{k}(v_{i-2k}-dw_{i-1-2k})=0$
Forcing $v_{i-2k}=dw_{i-1-2k}=\operatorname{\mathfrak{d}}w_{i-1-2k}+e\operatorname{\overline{\mathfrak{d}}}w_{i-1-2k}$ for all $k$. Since $v_{i}$ is harmonic, each $v_{i-2k}$ is symphonic, so we must have that $\operatorname{\mathfrak{d}}w_{i-1-2k}\in\operatorname{im}\operatorname{\mathfrak{d}}\cap\ker\operatorname{\overline{\mathfrak{d}}}$ and $\operatorname{\overline{\mathfrak{d}}}w_{i-1-2k}\in\operatorname{im}\operatorname{\overline{\mathfrak{d}}}\cap\ker\operatorname{\mathfrak{d}}$. By the $\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma then we must have that there are $\gamma_{i-2k},\rho_{i-2-2k}\in\operatorname{im}(\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}})$ (possibly zero) such that
$\operatorname{\mathfrak{d}}w_{i-1-2k}=\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\gamma_{i-2k}\qquad\operatorname{\overline{\mathfrak{d}}}w_{i-1-2k}=\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\rho_{i-2-2k}$
From here, we have that
$v_{i}$ $=\sum_{k\geq 0}e^{k}(\operatorname{\mathfrak{d}}w_{i-1-2k}+e\operatorname{\overline{\mathfrak{d}}}w_{i-1-2k})$
$=\sum_{k\geq 0}e^{k}(\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\gamma_{i-2k}+e\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\rho_{i-2(1+k)})$
$=\sum_{k\geq 0}e^{k}\left(d\delta\frac{\gamma_{i-2k}}{1-(i-2k)}+ed\delta\frac{\rho_{i-2(1+k)}}{1-i+2(1+k)}\right)$
$=d\delta\sum_{k\geq 0}e^{k}\left(\frac{\gamma_{i-2k}}{1-(i-2k)}+e\frac{\rho_{i-2(1+k)}}{1-i+2(1+k)}\right)$
Where we repeatedly use the $\mathfrak{g}$ relations, and make note that the constants that appear in the denominator are non-vanishing. The only possible point of confusion is why the $\gamma_{i-2k}$ and $\rho_{i-2(1+k)}$ are primitive. If they’re not, then taking a primitive decomposition of them, it’s easy to see that we can adjust by constants to force this to be true. A similar argument holds for the case when $v\in\operatorname{im}(\delta)$. ∎
Lemma 2.20.
The $\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma holds for $V\in\mathrm{Ob}(\mathscr{C}_{n})$ if the $d\delta$-lemma does
Proof.
Suppose that $v_{i}\in\ker(\operatorname{\mathfrak{d}})\cap\ker(\operatorname{\overline{\mathfrak{d}}})\cap\operatorname{\mathcal{P}}_{i}$ and recall that this forces $v_{i}$ to be harmonic and each of it’s primitive decomposition terms to be symphonic. First, suppose that $v_{i}\in\operatorname{im}(\operatorname{\overline{\mathfrak{d}}})$, i.e. $v_{i}=\operatorname{\overline{\mathfrak{d}}}\beta_{i+1}$. Then taking primitive decompositions we have
$v_{i}=\sum_{k\geq 0}e^{k}v_{i-2k}\qquad\beta_{i+1}\sum e^{k}\beta_{i+1-2k}$
Using the coexactness we have must have that
$v_{i-2k}=\operatorname{\overline{\mathfrak{d}}}\beta_{i+1-2k}=\delta\left(\frac{\beta_{i+1-2k}}{-i+2k}\right)$
So $v_{i-2k}\in\operatorname{im}(\delta)\cap\ker(d)=\operatorname{im}(d\delta)$. This then gives us that there exist $\rho_{i-2k}$ such that
$v_{i-2k}=d\delta\rho_{i-2k}=\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}(1-i+2k)\rho_{i-2k}$
So we have
$v_{i}=\sum_{k\geq 0}e^{k}v_{i-2k}=\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\sum_{k\geq 0}e^{k}((1-i+2k)\rho_{i-2k})$
Again it may be unclear as to why the $\rho_{i-2k}$ are primitive, but this can be chosen to happen as before. Now, suppose that $v_{i}\in\operatorname{im}(\operatorname{\mathfrak{d}})$ so $v_{i}=\operatorname{\mathfrak{d}}\beta_{i-1}$ for some $\beta_{i-1}$. Then equating primitive decompositions we have that $v_{i-2k}=\operatorname{\mathfrak{d}}\beta_{i-1-2k}$ where
$v_{i}=\sum_{k\geq 0}e^{k}v_{i-2k}\qquad\beta_{i-1}=\sum_{k\geq 0}e^{k}\beta_{i-1-2k}$
Since $v_{i}$ is symphonic, we have that $\operatorname{\overline{\mathfrak{d}}}\operatorname{\mathfrak{d}}\beta_{i-1-2k}=0$ for all $k$. We then have that
$d\delta\beta_{i-1}$ $=\sum d\delta e^{k}\beta_{i-1-2k}$
$=\sum_{k\geq 0}e^{k}d\delta\beta_{i-1-2k}$
$=\sum_{k\geq 0}\frac{1}{2(1-k)-i}e^{k}\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\beta_{i-1-2k}$
$=\sum_{k\geq 0}\frac{-1}{2(1-k)-i}e^{k}\operatorname{\overline{\mathfrak{d}}}\operatorname{\mathfrak{d}}\beta_{i-1-2k}$
$=0$
So $\delta d\beta_{i-1}=0$, meaning that $d\beta_{i-1}\in\ker(\delta)\cap\operatorname{im}(d)$. By the $d\delta$-lemma then we have that there exist $\gamma_{i}$ such that $d\beta_{i-1}=d\delta\gamma_{i}$. Taking a primitive decomposition of $\gamma_{i}$ we have that
$d\delta\gamma_{i}$ $=\sum_{k\geq 0}e^{k}d\delta\gamma_{i-2k}$
$=\sum_{k\geq 0}(1-i+2k)e^{k}\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}\gamma_{i-2k}$
$=\operatorname{\overline{\mathfrak{d}}}\operatorname{\mathfrak{d}}\sum_{k\geq 0}(i-1-2k)e^{k}\gamma_{i-2k}$
Which shows us that
$v_{i}=\operatorname{\mathfrak{d}}\beta_{i-1}-e\operatorname{\overline{\mathfrak{d}}}\beta_{i-1}=\operatorname{\overline{\mathfrak{d}}}\left(\operatorname{\mathfrak{d}}\sum_{k\geq 0}(i-1-2k)e^{k}\gamma_{i-2k}-e\beta_{i-1}\right)$
That is, $v_{i}\in\operatorname{im}(\operatorname{\overline{\mathfrak{d}}})$ from which the result follows from the previous case. ∎
Theorem 2.21.
Let $V\in\mathrm{Ob}(\mathscr{C}_{n})$. Then the following are equivalent
  1. (1)
    The $d\delta$-lemma holds for $V$
  2. (2)
    The map $[e^{i}]:H^{-i}(V,d)\rightarrow H^{i}(V,d)$ is an isomorphism for all $i$
  3. (3)
    The inclusion $(\ker(\delta),d)\hookrightarrow(V,d)$ is a quasi-isomorphism
  4. (4)
    The $\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma holds for $V$
Proof.
The first three are equivalent by 2.12. The first and last are equivalent by Lemmas 2.20 and 2.19
3. Lefschetz Algebroids
3.1. Symplectic Lie Algebroids
Lie Algebroids were originally introduced by Pradines [zbMATH03247872] as a generalization of the tangent bundle of a manifold to the Differentiable Groupoid setting.
Definition 17.
A “Lie Algebroid” $(\mathcal{A},\rho,M)$ over a $C^{\infty}$-manifold $M$ is a vector bundle $\mathcal{A}\rightarrow M$, equipped with an “anchor” $\rho:\mathcal{A}\rightarrow TM$, and a bracket $[\cdot,\cdot]_{\mathcal{A}}$ on sections of $\mathcal{A}$ that is bilinear, alternating, and satisfies the Jacobi identity, such that
  1. (1)
    $[X,fY]_{\mathcal{A}}=f[X,Y]_{\mathcal{A}}+(\rho(X)\cdot f)Y$ for all $X,Y\in\Gamma(\mathcal{A})$ and $f\in C^{\infty}(M)$
  2. (2)
    $\rho([X,Y]_{\mathcal{A}})=[\rho(X),\rho(Y)]_{TM}$ for all $X,Y\in\Gamma(\mathcal{A})$
Definition 18.
Given a Lie Algebroid $(\mathcal{A},\rho,M)$, we can define the “deRham-Chevalley-Eilenberg complex” of $\mathcal{A}$, where our cochain complex is made of
$\Omega^{k}_{\mathcal{A}}(M)=\Gamma(\Lambda^{k}\mathcal{A}^{*})$
And we define
$\Omega^{*}_{\mathcal{A}}(M)=\bigoplus_{k}\Omega^{k}_{\mathcal{A}}(M)$
This comes with an $\mathcal{A}$-differential, $d_{\mathcal{A}}:\Omega^{k}_{\mathcal{A}}(M)\rightarrow\Omega^{k+1}_{\mathcal{A}}(M)$, defined as
$d_{\mathcal{A}}\eta(X_{1},...,X_{k+1})$ $=\sum_{i=1}^{k+1}(-1)^{i+1}\rho(X_{i})\cdot\eta(X_{1},...,\widehat{X_{i}},...,X_{k+1})$
$+\sum_{i<j}(-1)^{i+j}\eta([X_{i},X_{j}],X_{1},...,\widehat{X_{i}},...,\widehat{X_{j}},...,X_{k+1})$
For $X_{i}\in\Gamma(\mathcal{A})$ and $\eta\in\Omega^{k}_{\mathcal{A}}(M)$.
A standard computation shows that $d_{\mathcal{A}}^{2}=0$, and as such we can consider the cohomology of this complex
Definition 19.
Let $(A,\rho,M)$ be a Lie Algebroid. The “Lie Algebroid Cohomology” of $\mathcal{A}$, denoted $H^{k}_{\mathcal{A}}(M)$ is the cohomology of the complex
[…]
that is
$H_{\mathcal{A}}^{k}(M)=\frac{\ker\left(\Omega^{k}_{\mathcal{A}}(M)\xrightarrow{d_{\mathcal{A}}}\Omega^{k+1}_{\mathcal{A}}(M)\right)}{\operatorname{im}\left(\Omega^{k-1}_{\mathcal{A}}(M)\xrightarrow{d_{\mathcal{A}}}\Omega^{k}_{\mathcal{A}}(M)\right)}$
which we can rephrase busing $\mathcal{A}$-closed and exact terminology
$H_{\mathcal{A}}^{k}(M)=\frac{d_{\mathcal{A}}\text{-closed forms}}{d_{\mathcal{A}}\text{-exact forms}}$
The standard Cartan Calculus from the Tangent Bundle also extends to Lie Algebroids, however we’ll only use the notion of the exterior and interior products
Definition 20.
Given $X\in\Gamma(\mathcal{A})$, we define “$\mathcal{A}$-interior multiplication” with $X$, or “contraction by $X$”, denoted $\iota_{X}$, as the operator $\iota_{X}:\Omega^{k}_{\mathcal{A}}(M)\rightarrow\Omega^{k-1}_{\mathcal{A}}(X)$ such that
$\iota_{X}\eta(X_{1},...,X_{k-1})=\eta(X,X_{1},...,X_{k-1})$
Where $X_{i}\in\Gamma(\mathcal{A})$ and $\eta\in\Omega_{\mathcal{A}}^{k}(M)$
Symplectic Lie Algebroids were originally introduced by Nest and Tsygan [nest1999deformations] and much work has been done on them in recent years by may authors. As in classical Symplectic Geometry, Symplectic Lie Algebroids arise naturally as the “phase space” of any Lie algebroid as originally uncovered by Martínez [Martinez]. Most recently, Lin et al. [Lin_2023] proved an analogue of the Darboux-Marsden-Weinstien Theorem and described the processes of symplectic reduction in this context.
Definition 21.
Let $(\mathcal{A},\rho,M)$ be a Lie Algebroid. An “$\mathcal{A}$-Symplectic form” $\omega$ is a $d_{\mathcal{A}}$-closed, nondegenerate element of $\Omega^{2}_{\mathcal{A}}(M)$
This allows us to define Symplectic Lie Algebroids
Definition 22.
A “Symplectic Lie Algebroid” is a Lie Algebroid $\mathcal{A}$ equipped with a $\mathcal{A}$-symplectic form.
Corollary 3.0.1.
Let $(\mathcal{A}\rightarrow M,\rho,\omega)$ be a Symplectic Lie Algebroid. Then the fiber dimension of $\mathcal{A}$ is even and $\mathcal{A}$ is orientable
Proposition 3.1 (Martinez).
Let $B\rightarrow N$ be an arbitrary Lie Algebroid, and $\pi:B^{*}\rightarrow N$ be the projection. Define the bundle
$\mathcal{A}=\pi^{!}B=TB^{*}\times_{TN}B$
Which is a Lie Algebroid over $B^{*}$. Then $\omega=-d_{\mathcal{A}}\alpha$ is a canonical symplectic structure on $B^{*}$, where $\alpha\in\Omega_{\mathcal{A}}^{1}(B^{*})$ is defined as follows after identifying elements of $\mathcal{A}$ as tuples in the product bundle
$\alpha(x,p,v,b)=p(b)\qquad x\in N,p\in B^{*}_{x},v\in T_{p}B^{*},b\in B_{x}$
Satisfying $T_{p}\pi(v)=\rho_{B}(b)$.
Proof.
See originally Martinez [Martinez] or Lin [Lin_2023]
Proposition 3.2.
Given Symplectic Lie Algebroids $(\mathcal{A}_{1}\rightarrow M_{1},\omega_{1})$ and $(\mathcal{A}_{2}\rightarrow M_{2},\omega_{2})$, the product $(A_{1}\times A_{2}\rightarrow M_{1}\times M_{2},\omega)$ where $\omega=\pi_{1}^{*}\omega_{1}+\pi_{2}^{*}\omega_{2}$ is a Symplectic Lie Algebroid
Proof.
The product of two Lie Algebroids has a unique Lie Algebroid structure [GroupoidAlgebroidNotes], so we need to check that $\omega$ is a symplectic form. It’s clear that it’s nondegenerate as it’s made from the pullback of two nondegenerate forms. To see that it’s closed note that $\pi_{i}$ are Lie Algebroid morphisms, and a map $f:\mathcal{A}\rightarrow\mathcal{B}$ is a Lie Algebroid morphism if and only if the induced map $f^{*}:\Omega^{*}_{\mathcal{B}}\rightarrow\Omega^{*}_{\mathcal{A}}$ is a chain map. Using this we have that $\pi_{i}^{*}$ are chain maps and so
$d_{\mathcal{A}_{1}\times\mathcal{A}_{2}}\omega=d_{\mathcal{A}_{1}\times\mathcal{A}_{2}}\pi_{1}^{*}\omega_{1}+d_{\mathcal{A}_{1}\times\mathcal{A}_{2}}\pi_{2}^{*}\omega_{2}=\pi_{1}^{*}(d_{\mathcal{A}_{1}}\omega_{1})+\pi_{2}^{*}(d_{\mathcal{A}_{2}}\omega_{2})=0$
Lemma 3.1.
Let $(\mathcal{A}\rightarrow M,\rho,\omega)$ be a Symplectic Lie Algebroid of fiber dimension $2m$. Then for all $p\in M$, there exists a smooth symplectic local frame of $\mathcal{A}$, i.e. there exists a frame $\{u_{i},v_{i}\}_{i=1}^{m}\subset\Gamma(\mathcal{A})$ such that $\{u_{i}(q),v_{i}(q)\}_{i=1}^{m}$ is a symplectic basis of $\mathcal{A}_{q}$ for all $q\in U$, an open neighborhood of $p$.
Proof.
Fix $p\in M$, and let $B_{0}=\{s_{i}\}_{i=1}^{2m}$ be a smooth local frame of $\mathcal{A}$ over a neighborhood $V$ of $p$. Now, define $u_{1}=s_{1}$, and choose $\epsilon_{1}\in\Gamma(\mathcal{A}|_{U})$ such that
$\omega\left(u_{1},\frac{s_{2}+\epsilon_{1}}{C_{1}}\right)=1$
Which (after possibly restricting to a smaller open set $p\in V_{1}\subset V$) exists by non-degeneracy of $\omega$, and we rescale by $C_{1}$ to have the desired constant $1$, so that for $v_{1}:=\frac{s_{2}+\epsilon_{1}}{C_{1}}$ we have $\omega(u_{1},v_{1})=1$. Now, let
$u_{2}=s_{3}-\omega(s_{3},v_{1})u_{1}+\omega(s_{3},u_{1})v_{1}$
Then $\omega(u_{1},u_{2})=\omega(v_{1},u_{2})=0$ by design. For $v_{2}$, start by choosing $\epsilon_{2}\in\Gamma(\mathcal{A}|_{U})$ such that
$\omega\left(u_{2},\frac{s_{4}+\epsilon_{2}}{C_{2}}\right)=1$
Which again is possible by non-degeneracy of $\omega$ after possibly shrinking to $p\in V_{2}\subset V_{1}$. Call $v_{2}^{\prime}=\frac{s_{4}+\epsilon_{2}}{C_{2}}$, and now define
$v_{2}=v_{2}^{\prime}-\omega(v_{2}^{\prime},v_{1})u_{1}+\omega(v_{2}^{\prime},u_{1})v_{1}$
Then we have that
$\omega(u_{1},v_{2})=\omega(v_{1},v_{2})=0\qquad\omega(u_{2},v_{2})=1$
We can continue this process $m$ times giving us a frame of the desired form over $V_{m}:=U$. ∎
Note here that we’re not insisting that the algebroid $1$-forms that make up our local coframe of $\mathcal{A}^{*}$ are differentials of coordinate functions, or that they’re exact. In fact the coframe we define is given by
$\xi_{i}=u_{i}^{\flat}\qquad\mu_{i}=-v_{i}^{\flat}$
Where $\flat:\Gamma(\mathcal{A})\rightarrow\Gamma(\mathcal{A}^{*})$ is given by $X\mapsto\omega(X,-)$. The local frame $\{\mu_{i},\xi_{i}\}_{i=1}^{m}$ of $\mathcal{A}^{*}$ is dual to $\{u_{i},v_{i}\}_{i=1}^{m}$, but a geometrically inclined choice of normal coordinates that correspond to this are harder to pin down. There are cohomological obstructions to making such a choice as shown by Miranda and Scott [Miranda_2020] (Theorem 29). For this reason we will not make an appeal to a full “Darboux Theorem” though work has been done on this see [Lin_2023]. In many settings such as $b$-manifolds there are results on local normal forms, and in these settings the following sections results can be more easily understood geometrically, see [Matveeva_2023], [Miranda_2020].
3.2. Kahler-Weil Identities
Throughout this section, let $(\mathcal{A}\rightarrow M,\rho,\omega)$ be a Symplectic Lie Algebroid of fiber dimension $2m$. We introduce the main two operators involved in this result. The “Lefschetz” map
$L:\Omega^{k}_{\mathcal{A}}(M)$ $\rightarrow\Omega^{k+2}_{\mathcal{A}}(M)$
$\alpha$ $\mapsto\omega\wedge\alpha$
And if we let $\lambda$ denote the dual bivector to $\omega$, we define
$\Lambda:\Omega^{k}_{\mathcal{A}}(M)$ $\rightarrow\Omega^{k-2}_{\mathcal{A}}(M)$
$\alpha$ $\mapsto\iota_{\lambda}\alpha$
By our weaker analogue of the Darboux Theorem, Lemma 3.1 , we have local symplectic frames of $\mathcal{A}$ and $\mathcal{A}^{*}$ on an open set $U$ given by
$\mathcal{A}|_{U}=\mathrm{span}_{C^{\infty}(M)}\{u_{i},v_{i}\}_{i=1}^{m}\qquad\mathcal{A}^{*}|_{U}=\mathrm{span}_{C^{\infty}(M)}\{\mu_{i},\xi_{i}\}_{i=1}^{m}$
In such a frame we can write local expressions for $L$ and $\Lambda$
$L=\sum_{i}e_{\mu_{i}}e_{\xi_{i}}\qquad\Lambda=\sum_{i}\iota_{v_{i}}\iota_{u_{i}}$
In order to understand these operators better define $\star:\Omega^{k}_{\mathcal{A}}(M)\rightarrow\Omega^{2m-k}_{\mathcal{A}}(M)$ as the operator
$\alpha\wedge\star\beta=\lambda^{k}(\alpha,\beta)\frac{\omega^{m}}{m!}$
For all $\alpha,\beta\in\Omega_{\mathcal{A}}^{k}(M)$. A standard result shows the following
Proposition 3.3.
$\star\star=\mathrm{Id}$
Definition 23.
The “Brylinski” differential, $d_{\mathcal{A}}^{\star}:\Omega^{k}_{\mathcal{A}}(M)\rightarrow\Omega_{\mathcal{A}}^{k-1}(M)$ defined as follows for $\alpha\in\Omega^{k}_{\mathcal{A}}(M)$
$d_{\mathcal{A}}^{\star}\alpha=(-1)^{k+1}\star d_{\mathcal{A}}\star\alpha$
This operator also creates a chain complex on the $\Omega_{\mathcal{A}}^{\bullet}(M)$ in the opposite direction of $d_{\mathcal{A}}$ since $d_{\mathcal{A}}^{\star}\circ d_{\mathcal{A}}^{\star}=0$. The homology of this complex will not be explored here. As in the classical setting, we have the following definition
Definition 24.
We say that a form $\alpha\in\Omega^{*}_{\mathcal{A}}(M)$ is “$\mathcal{A}$-symplectic harmonic” or just “harmonic” if $d_{\mathcal{A}}\alpha=d_{\mathcal{A}}^{\star}\alpha=0$, and denote the set of harmonic $k$-forms as $\widehat{\Omega}_{\mathcal{A}}^{k}(M)$ and the set of all harmonic forms as $\widehat{\Omega}_{\mathcal{A}}^{*}(M)$.
Proposition 3.4.
$\Lambda=\star L\star$
Proof.
This is a strictly Linear Algebraic fact, see [Shlomonotes] for details ∎
Lemma 3.2.
If $\alpha\in\Omega^{k}_{\mathcal{A}}(M)$, then $[L,\Lambda]\alpha=(k-m)\alpha$
Proof.
These are local, $C^{\infty}(M)$-linear operators, so fix a point $p\in M$ and an open neighborhood $U\ni p$ such that we have the symplectic frames $\{u_{i},v_{i}\}_{i=1}^{m}$ of $\mathcal{A}$ and $\{\mu_{i},\xi_{i}\}_{i=1}^{m}$ of $\mathcal{A}^{*}$. From here we can proceed as follows in [Shlomonotes] on page 8. ∎
We then extend this result to the entire space of forms using the following definition
Definition 25.
We define the “counting operator” $H:\Omega^{*}_{\mathcal{A}}(M)\rightarrow\Omega^{*}_{\mathcal{A}}(M)$ as the map
$H=\sum_{k}(k-m)\pi_{k}$
Where $\pi_{k}:\Omega^{*}_{\mathcal{A}}(M)\rightarrow\Omega^{k}_{\mathcal{A}}(M)$ is the canonical projection
Proposition 3.5.
Given the operators $L,\Lambda$, and $H$, we have the following $\mathfrak{sl}_{2}$-commutation relations
$[L,\Lambda]=H\quad[H,L]=2L\quad[H,\Lambda]=-2\Lambda$
Proof.
The first identity is the result of the lemma 3.2, so let’s check the second one on $\alpha\in\Omega^{k}_{\mathcal{A}}(M)$
$[H,L]\alpha$ $=H(L\alpha)-L((k-m)\alpha)$
$=(k+2-m)L\alpha-(k-m)L\alpha$
$=2L\alpha$
The third identity follows by a similar argument
$[H,\Lambda]\alpha$ $=H(\Lambda\alpha)-\Lambda(H\alpha)$
$=(k-2-m)\Lambda\alpha-(k-m)\Lambda\alpha$
$=-2\Lambda\alpha$
Lemma 3.3.
We have the commutation relations
  1. (1)
    $[L,d_{\mathcal{A}}]=0$
  2. (2)
    $[H,d_{\mathcal{A}}]=d_{A}$
  3. (3)
    $[H,d_{\mathcal{A}}^{*}]=-d_{\mathcal{A}}^{*}$
  4. (4)
    $[\Lambda,d_{\mathcal{A}}^{*}]=0$
  5. (5)
    $[L,d_{\mathcal{A}}^{*}]=d_{\mathcal{A}}$
  6. (6)
    $[\Lambda,d_{\mathcal{A}}]=d_{\mathcal{A}}^{*}$
Proof.
We prove these in order for $\alpha\in\Omega_{\mathcal{A}}^{k}(M)$
  1. (1)
    Using that $\omega$ is closed we have
    $[L,d_{\mathcal{A}}]\alpha$ $=\omega\wedge d_{\mathcal{A}}\alpha-d_{\mathcal{A}}(\omega\wedge\alpha)$
    $=\omega\wedge d_{\mathcal{A}}\alpha-(d_{\mathcal{A}}\omega\wedge\alpha+\omega\wedge d_{\mathcal{A}}\alpha)$
    $=0$
  2. (2)
    Using only degree considerations we have
    $[H,d_{\mathcal{A}}]\alpha=Hd_{\mathcal{A}}\alpha-d_{\mathcal{A}}H\alpha=[(k+1)-n]d_{\mathcal{A}}\alpha-(k-n)d_{\mathcal{A}}\alpha=d_{\mathcal{A}}\alpha$
  3. (3)
    Again using only degree considerations we have that
    $[H,d_{\mathcal{A}}^{\star}]=Hd_{\mathcal{A}}^{\star}\alpha-d_{\mathcal{A}}^{\star}H\alpha=[(k-1)-n]d_{\mathcal{A}}^{\star}\alpha-(k-n)d_{\mathcal{A}}^{\star}\alpha=-d_{\mathcal{A}}^{\star}\alpha$
  4. (4)
    Using the star form of the operators we have
    $[\Lambda,d_{\mathcal{A}}^{\star}]\alpha=(-1)^{k+1}\star[L,d]\star\alpha=0$
  5. (5)
    Using the Jacobi identity we have
    $[L,d_{\mathcal{A}}^{\star}]+[d_{\mathcal{A}},[L,\Lambda]]+[\Lambda,[d_{\mathcal{A}},L]]$ $=0$
    Now we know the last term on the right hand side vanishes as $d_{\mathcal{A}}$ and $L$ commute, and using the Kähler-Weil identity we have
    $[L,d_{\mathcal{A}}^{\star}]=[H,d_{\mathcal{A}}]=d_{\mathcal{A}}$
  6. (6)
    Using the Jacobi identity along with the previous work we have that
    $[\Lambda,[L,d_{\mathcal{A}}^{\star}]]+[d_{\mathcal{A}}^{\star},[\Lambda,L]]+[L,[d_{\mathcal{A}}^{\star},\Lambda]]$ $=0$
    $[\Lambda,d]+[d_{\mathcal{A}}^{\star},-H]+0$ $=0$
    Giving us that
    $[\Lambda,d]=d_{\mathcal{A}}^{\star}$
Proposition 3.6.
The space of $\mathcal{A}$-forms of a Symplectic Lie Algebroid of fiber dimension $2m$ with the operators $L,\Lambda,H,d_{\mathcal{A}},d_{\mathcal{A}}^{\star}$ is an object of $\mathscr{C}_{m}$ under the assignment
$e\mapsto L\quad f\mapsto\Lambda\quad h\mapsto H\quad d\mapsto d_{\mathcal{A}}\quad\delta\mapsto d_{\mathcal{A}}^{\star}$
Proof.
The commutation relations follow proposition 3.5 and lemma 3.3
Theorem 3.4.
Suppose $(\mathcal{A}\rightarrow M,\rho,\omega)$ is a Symplectic Lie Algebroid of fiber dimension $2m$. Then the following statements are equivalent:
  1. (1)
    The inclusion $(\ker(d_{\mathcal{A}}^{\star}),d_{\mathcal{A}})\hookrightarrow(\Omega_{\mathcal{A}}^{*}(M),d)$ is a quasi-isomorphism
  2. (2)
    For all $0\leq k\leq m$, the map $[L]^{k}:H_{\mathcal{A}}^{m-k}(M)\rightarrow H_{\mathcal{A}}^{m+k}(M)$ is an isomorphism
  3. (3)
    $(\mathcal{A}\rightarrow M,\rho,\omega)$ satisfies the $\mathcal{A}$-$dd^{\star}$-lemma
  4. (4)
    $(\mathcal{A}\rightarrow M,\rho,\omega)$ satisfies the $\mathcal{A}$-$\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma
where the first three statements hold equivalence weakly, i.e. in degree as in the statement of Theorem 2.14
Proof.
This is the statement of Theorem 2.21 under the representation ∎
Definition 26.
We say that a Symplectic Lie Algebroid $(\mathcal{A},\rho,\omega)$ is a “Lefschetz Algebroid” if any of the four conditions in the preceding theorem hold
Corollary 3.4.1.
For any Lefschetz Algebroid $(\mathcal{A}\rightarrow M,\rho,\omega)$, the space $H^{*}_{\mathcal{A}}(M)$ is an $\mathfrak{sl}_{2}(\mathbb{R})$-module
Proof.
This follows from Corollary 2.12.1
Corollary 3.4.2.
The deRham-Chevalley-Eilenberg complex of a Lefschetz Lie Algebroid is formal
Proof.
Condition 1 in Theorem 3.4 shows that a Lefschetz Lie Algebroid satisfies the criteria for the in Manin [manin1998constructions], from which the result follows. ∎
We also have the weak version of this theorem
Theorem 3.5.
Suppose $(\mathcal{A}\rightarrow M,\rho,\omega)$ is a Symplectic Lie Algebroid of fiber dimension $2m$, and $0\leq s\leq m$. Then the following statements are equivalent:
  1. (1)
    The inclusion $(\ker(d_{\mathcal{A}}^{\star}),d_{\mathcal{A}})\hookrightarrow(\Omega_{\mathcal{A}}^{\bullet}(M),d)$ is a quasi-isomorphism in degrees $\leq s$ or $\geq s+m$
  2. (2)
    The map $[L]^{k}:H_{\mathcal{A}}^{m-k}(M)\rightarrow H_{\mathcal{A}}^{m+k}(M)$ is an isomorphism for $k\geq s$
  3. (3)
    The $\mathcal{A}$-$d_{\mathcal{A}}d_{\mathcal{A}}^{\star}$-lemma holds up to degree $s$
Proof.
This is the statement of Theorem 2.14 under the representation. ∎
3.3. Cohomology with Coefficients
This whole story works as well for Lie Algebroid cohomology with coefficients in a representation of the Lie Algebroid. This extension was originally pointed out in [YAN1996143] for traditional vector bundles with connections.
Definition 27.
Let $\mathcal{A}\rightarrow M$ be a Lie Algebroid over a smooth manifold $M$. An “$\mathcal{A}$-connection” on a vector bundle $E\rightarrow M$ is an $\mathbb{R}$-bilinear map
${}^{\mathcal{A}}\nabla:\Gamma\mathcal{A}\times\Gamma E$ $\rightarrow\Gamma E$
$(X,s)$ $\mapsto\nabla_{X}s$
Which is $C^{\infty}(M)$ linear in the first entry and satisfies
${}^{\mathcal{A}}\nabla_{X}(fs)=f{}^{\mathcal{A}}\nabla_{X}s+\mathcal{L}_{\rho(X)}(f)s$
The “curvature” of the connection ${}^{\mathcal{A}}\nabla$ is the $\operatorname{End}(E)$-valued algebroid $2$-form
$R_{\nabla}(X,Y)={}^{\mathcal{A}}\nabla_{X}{}^{\mathcal{A}}\nabla_{Y}-{}^{\mathcal{A}}\nabla_{Y}{}^{\mathcal{A}}\nabla_{X}-{}^{\mathcal{A}}\nabla_{[X,Y]}$
We say that ${}^{\mathcal{A}}\nabla$ is “flat” if $R_{\nabla}\equiv 0$
Definition 28.
Let $\mathcal{A}\rightarrow M$ be a Lie Algebroid over a smooth manifold. A “representation” of $\mathcal{A}$ is a vector bundle $E$ equipped with a flat $\mathcal{A}$-connection.
Using the connection we can create the space of forms with coefficients in a representation which inherits a grading from $\Omega^{k}_{\mathcal{A}}(M)$ and comes with a differential given by the covariant derivative
Definition 29.
Let $\mathcal{A}\rightarrow M$ be a Lie Algebroid and $E$ be a representation of $\mathcal{A}$. Then the space of “Algebroid $k$-forms with coefficients in $E$” is defined as
$\Omega_{\mathcal{A}}^{k}(M,E)=\Gamma(\bigwedge^{k}\mathcal{A}^{*}\otimes E)$
We can assemble these into the differential graded module of all forms
$\Omega_{\mathcal{A}}^{*}(M,E)=\bigoplus_{k\geq 0}\Omega^{k}_{\mathcal{A}}(M,E)$
With differential ${}^{\mathcal{A}}\nabla:\Omega_{\mathcal{A}}^{k}(M,E)\rightarrow\Omega_{\mathcal{A}}^{k+1}(M,E)$ given by the induced covariant derivative, i.e. we extend ${}^{\mathcal{A}}\nabla:\Omega_{\mathcal{A}}^{0}(M,E)\rightarrow\Omega_{\mathcal{A}}^{1}(M,E)$ via enforcing
${}^{\mathcal{A}}\nabla(\alpha\otimes s):=(d_{\mathcal{A}}\alpha)\otimes s+(-1)^{|\alpha|}\alpha\otimes({}^{\mathcal{A}}\nabla s)$
For $\alpha\otimes s\in\Omega^{k}_{A}(M,E)$. Since the connection is flat, this is a cochain complex.
Proposition 3.7.
Let $(\mathcal{A},\rho,\omega)$ be a Symplectic Lie Algebroid of fiber dimension $2m$ and $E$ be a representation of $\mathcal{A}$. Then we have operators
$L_{\nabla}:\Omega^{k}_{\mathcal{A}}(M,E)$ $\rightarrow\Omega_{\mathcal{A}}^{k+2}(M,E)$
$\alpha\otimes e$ $\mapsto(\omega\wedge\alpha)\otimes e=(L\alpha)\otimes e$
$\Lambda_{\nabla}:\Omega_{\mathcal{A}}^{k}(M,E)$ $\rightarrow\Omega_{\mathcal{A}}^{k-2}(M,E)$
$\alpha\otimes e$ $\mapsto(\iota_{\omega^{-1}}\alpha)\otimes e=(\Lambda\alpha)\otimes e$
$H_{\nabla}:\Omega_{\mathcal{A}}^{k}(M,E)$ $\rightarrow\Omega_{\mathcal{A}}^{k}(M,E)$
$\alpha\otimes e$ $\mapsto((k-m)\alpha)\otimes e=(H\alpha)\otimes e$
${}^{\mathcal{A}}\nabla^{*}:\Omega_{\mathcal{A}}^{k}(M,E)$ $\rightarrow\Omega_{\mathcal{A}}^{k-1}(M,E)$
$\alpha\otimes e$ $\mapsto[\Lambda_{\nabla},{}^{\mathcal{A}}\nabla](\alpha\otimes e)$
These together with ${}^{\mathcal{A}}\nabla$ satisfy the relations of $\mathfrak{g}$, and as such the space of forms with coefficients are an object of $\mathscr{C}_{n}$.
Proof.
The proof follows similarly to that of Lemma 3.3
Definition 30.
Let $(\mathcal{A},\rho,M)$ be a Symplectic Lie Algebroid of fiber dimension $2m$, $E\rightarrow M$ be a representation of $\mathcal{A}$ with connection ${}^{\mathcal{A}}\nabla$, and $\alpha\in\Omega^{*}_{\mathcal{A}}(M,E)$. We say that $\alpha$ is “${}^{\mathcal{A}}\nabla$-Symplectic Harmonic” if ${}^{\mathcal{A}}\nabla\alpha={}^{\mathcal{A}}\nabla^{*}\alpha=0$.
Corollary 3.5.1.
Let $(\mathcal{A}\rightarrow M,\rho,\omega)$ be a Symplectic Lie Algebroid of fiber dimension $2m$, with representation $(E,\nabla)$. Then the following are equivalent:
  1. (1)
    The map $[L]^{k}:H^{m-k}_{\mathcal{A}}(M,E)\rightarrow H^{m+k}_{\mathcal{A}}(M,E)$ given by $[\omega]^{k}\wedge-$ is an isomorphism for all $k$
  2. (2)
    The inclusion of complexes $(\ker({}^{\mathcal{A}}\nabla^{*}),{}^{\mathcal{A}}\nabla)\hookrightarrow(\Omega_{\mathcal{A}}^{\bullet}(M,E),{}^{\mathcal{A}}\nabla)$ is a quasi-isomorphism
  3. (3)
    The $(\mathcal{A},E)$-$d\delta$-lemma holds for $\mathcal{A}$
  4. (4)
    The $(\mathcal{A},E)$-$\operatorname{\mathfrak{d}}\operatorname{\overline{\mathfrak{d}}}$-lemma holds for $\mathcal{A}$
and the first three statements hold equivalence weakly, i.e. in degree as in the statement of Theorem 2.14.
The above corollary is the statement of Theorems 2.14 and 2.21. Taking $E$ to be the trivial line bundle recovers Theorem 3.4, however it would be interesting to look at what this theorem means for the choice
$E=Q_{\mathcal{A}}=\bigwedge^{\text{top}}\mathcal{A}^{*}\otimes\bigwedge^{\text{top}}T^{*}M$
As defined in [evens1996transversemeasuresmodularclass] as this always admits the structure of a representation of $\mathcal{A}$.
4. Examples
4.1. Kodaria-Thurston Manifold
One of the first examples of a symplectic manifold that does not admit a Kähler structure is given by the “Kodaira-Thurston” manifold $KT^{4}$ which is constructed as follows. Take $\mathbb{R}^{4}$ with coordinates $(x_{1},x_{2},x_{3},x_{4})$ and choose $a,b,c,d\in\mathbb{Z}$. Then consider the quotient of $\mathbb{R}^{4}$ by identifying
$(x_{1},x_{2},x_{3},x_{4})\sim(x_{1}+a,x_{2}+b,x_{3}+c,x_{4}+d-bx_{3})$
The resulting quotient is the connected smooth manifold $KT^{4}$. We can see that we have a basis of one forms $\{e_{i}\}_{i=1}^{4}$ given by
$e_{1}=dx_{1}\qquad e_{2}=dx_{2}\qquad e_{3}=dx_{3}\qquad e_{4}=dx_{4}+x_{2}dx_{3}$
We have a symplectic structure on $KT^{4}$ given by $\omega=e_{1}\wedge e_{2}+e_{3}\wedge e_{4}$, and the deRham cohomology of $KT^{4}$ has given by generators
$H_{dR}^{0}(KT^{4})$ $=\langle[1]\rangle$ $H_{dR}^{1}(KT^{4})$ $=\langle[e_{1}],[e_{2}],[e_{3}]\rangle$
$H_{dR}^{2}(KT^{4})$ $=\langle\omega,[e_{12}-e_{34}],[e_{13}],[e_{24}]\rangle$ $H_{dR}^{3}(KT^{4})$ $=\langle[\omega\wedge e_{1}],[\omega\wedge e_{2}],[\omega\wedge e_{4}]\rangle$
$H_{dR}^{4}(KT^{4})$ $=\langle[\frac{1}{2}\omega^{2}]\rangle$
Now we can consider the Lefschetz maps
$[L]:H_{dR}^{1}(KT^{4})\rightarrow H_{dR}^{3}(KT^{4})\qquad[L]^{2}:H_{dR}^{0}(KT^{4})\rightarrow H_{dR}^{4}(KT^{4})$
The map $[L]^{2}$ is clearly an isomorphism, but $[L]$ is not as it has nontrivial kernel, and doesn’t surject either. This satisfies the criteria of Theorem 3.4 of degree $s=1$ where we take $\mathcal{A}=TM$, the identity as the anchor, and the differential as the classical deRham differential.
4.2. Symplectic Lie Algebras of Dimension Four
If we consider Symplectic Lie Algebras as Symplectic Lie Algebroids over the one-point space, the Lie Algebroid Cohomology is exactly the Chevalley-Eilenberg cohomology, and our theorem can be checked by examining the surjectivity of the maps $[L]^{k}:H^{m-k}(\mathfrak{g})\rightarrow H^{m+k}(\mathfrak{g})$. For clarification sake by a Symplectic Lie Algebra we mean a (real) Lie Algebra $\mathfrak{g}$ equipped with a nondegenerate $\omega\in\bigwedge^{2}\mathfrak{g}^{*}$ that is closed under the Chevalley-Elienberg differential. The 2-dimensional case has only two isomorphism classes; the abelian one which satisfies Theorem 3.5 and the other one which does not. Things get more interesting in the 4-dimensional case, where the number of symplectomorphism classes grows significantly. In these case we need to check the following two maps in cohomology in the diagram:
00{0}${{H^{0}(\mathfrak{g})}}$${{H^{1}(\mathfrak{g})}}$${{H^{2}(\mathfrak{g})}}$${{H^{3}(\mathfrak{g})}}$${{H^{4}(\mathfrak{g})}}$00{0}$\scriptstyle{[L]^{1}}$$\scriptstyle{[L]^{2}}$
A result by Ovando [ovando2004dimensional] gives a classification of all 4-dimensional Real Solvable Symplectic Lie Algebras up to symplectomorphism. Recomputing the cohomology of these classes we can check by hand which maps $[L]^{i}$ are isomorphisms, surjections, or neither. Using the notation of Ovando and numbering for different symplectic forms, we can check the maps in cohomology which results in the following table where $s$ denotes the $s$ as in Theorem 3.5, i.e. $s=2$ is the Lefschetz condition.
Lie Algebra $s$ $[L]$ $[L]^{2}$
$\mathfrak{rh}_{3}$ 1 Neither Isomorphism
$\mathfrak{rr}_{3,0}$ 0 Surjection Surjection
$\mathfrak{rr}_{3,-1}$ 2 Isomorphism Isomorphism
$\mathfrak{rr}_{3,0}^{\prime}$ 2 Isomorphism Isomorphism
$\mathfrak{r}_{2}\mathfrak{r}_{2}$ 0 Surjection Surjection
$\mathfrak{r}_{2}^{\prime}$ 1 0 Isomorphism
$\mathfrak{n}_{4}$ 1 0 Isomorphism
$\mathfrak{r}_{4,0}$ 0 Surjection Surjection
$\mathfrak{r}_{4,-1}$ 0 0 Surjection
$\mathfrak{r}_{4,-1,\beta},\beta\in(-1,0)$ 0 Isomorphism Surjection
$\mathfrak{r}_{4,-1,-1}$ 0 Neither Surjection
$\mathfrak{r}_{4,\alpha,-\alpha},\alpha\in(-1,0)$ 0 Surjection Surjection
$\mathfrak{r}_{4,0,\delta}^{\prime},\delta>0$ 0 Isomorphism Surjection
$\mathfrak{d}_{4,1}$ 0 Surjection Surjection
$\mathfrak{d}_{4,2}$, $\omega_{1}$ 0 0 Surjection
$\mathfrak{d}_{4,2}$, $\omega_{2}$ 0 Isomorphism Surjection
$\mathfrak{d}_{4,2}$, $\omega_{3}$ 1 Isomorphism Surjection
$\mathfrak{d}_{4,\lambda},\lambda\geq 1/2$ 0 Surjection Surjection
$\mathfrak{d}_{4,\delta}^{\prime},\delta>0$ 0 Surjection Surjection
$\mathfrak{h}_{4}$ 0 Surjection Surjection
Table 1. Lie Algebra Lefschetz Maps
4.3. Six-Dimensional Nilmanifolds
Nilmanifolds are homogeneous spaces $N/\Gamma$ where $N$ is a simply connected nilpotent real Lie Group, and $\Gamma$ is a lattice in $N$ of maximal rank. Such spaces have been classified up to isomorphism in dimension $6$, and there are $34$ such isomorphism classes. It is known [NilpLieAlg] that exactly 26 of the 34 isomorphism classes admit symplectic structures. The data of such a space is typically presented by two pieces of information: Its class, given as a $6$-tuple, and its symplectic form. For example, the tuple $(0,0,0,0,12,14+25)$ with symplectic form $13+26+45$ conveys the data of the Lie Algebra with dual generators $\{e^{i}\}_{i=1}^{6}$ such that
$de^{1}=de^{2}$ $=de^{3}=de^{4}=0$
$de^{5}$ $=e^{1}\wedge e^{2}$
$de^{6}$ $=e^{1}\wedge e^{4}+e^{2}\wedge e^{5}$
with symplectic form $\omega=e^{1}\wedge e^{3}+e^{2}\wedge e^{6}+e^{4}\wedge e^{5}$. Only one of these classes admit a Kähler structure, and the other classes have provided interesting (counter)-examples in Symplectic Geometry. The fact that the Lie Algebra data is sufficient to classify these is evident in the following theorem:
Theorem (Nomizu [MR64057]).
The deRham complex $\Omega^{\bullet}(M)$ of a nilmanifold $M=N/\Gamma$ is quasi-isomorphic to the complex $\bigwedge^{\bullet}\mathfrak{n}^{*}$ of left-invariant forms on $N$, thus $H_{dR}^{\bullet}(M)\cong H_{CE}^{\bullet}(\mathfrak{n})$.
Examining the maps $[L]^{k}:H^{3-k}(\mathfrak{n})\rightarrow H^{3+k}(\mathfrak{n})$ where $[\alpha]\mapsto[\omega^{k}\wedge\alpha]$ for $k=1,2,3$ allows to check when Theorem 3.5 holds. The first of these maps $[L]^{3}$ is assured to be an isomorphism by compactness of the nilmanifold, and the remaining two are explicitly calculable given the standard data of a symplectic nilmanifold described above. By Poincareé Duality, the maps $[L]^{k}$ are between vector spaces of the same finite dimension, and so being surjective is equivalent to being an isomorphism. Only one class will admit all three isomorphisms, as in this setting all three maps being isomorphisms is equivalent to admitting Kähler structure [MR993739]. Table 2 lists whether or not a given nilmanifold class has maps $[L]^{k}$ that are isomorphisms; Yes indicating it’s an isomorphism, and No indicating it’s not an isomorphism.
Signature Symplectic Form $[L]$ $[L]^{2}$ $[L]^{3}$
(0,0,12,13,14,15) 16+34-25 No No Yes
(0,0,12,13,14,23+15) 16+24+34-26 No No Yes
(0,0,12,13,23,14) 15+24+34-26 Yes No Yes
(0,0,12,13,23,14-25) 15+24-35+16 Yes No Yes
(0,0,12,13,23,14+25) 15+24+35+16 Yes No Yes
(0,0,12,13,14+23,24+15) $16+2\times 34-25$ No No Yes
(0,0,0,12,13,14+23) $16-2\times 34-25$ No No Yes
(0,0,0,12,13,24) 26+14+35 No No Yes
(0,0,0,12,13,14) 16+24+35 No No Yes
(0,0,0,12,13,23) 15+24+36 No No Yes
(0,0,0,12,14,15+23) 13+26-45 No No Yes
(0,0,0,12,14+15+23+24) 13+26-45 No No Yes
(0,0,0,12,14,15+24) 13+26-45 No No Yes
(0,0,0,12,14,15) 13+26-45 No No Yes
(0,0,0,12,14,13+42) 15+26+34 No No Yes
(0,0,0,12,14,23+24) 16-34+25 No No Yes
(0,0,0,12,14,15+34) 16+35+24 No No Yes
(0,0,0,12,14+23,13+42) $15+2\times 26+34$ No No Yes
(0,0,0,0,12,15) 16+25+34 No No Yes
(0,0,0,0,12,14+25) 13+26+45 No No Yes
(0,0,0,0,12,14+23) 3+26+45 No No Yes
(0,0,0,0,12,34) 15+36+24 No No Yes
(0,0,0,0,12,13) 16+25+34 No No Yes
(0,0,0,0,13+42, 14+23) 16+25+34 No No Yes
(0,0,0,0,0,12) 16+23+45 No No Yes
(0,0,0,0,0,0) 12+34+56 Yes Yes Yes
Table 2. Six-Dimensional Nilmanifold Lefschetz Maps
4.4. $E$-manifolds
$E$-manifolds are a generalization of a collection of different geometric settings: $b$-manifolds, $b^{m}$-manifolds, $c$-manifolds, and Regular Foliations. Miranda and Scott have thoroughly explored the cohomology and symplectic geometry of $E$-manifolds in many settings, e.g. [Miranda_2020]. The definition given in the above paper of Miranda and Scott of an $E$-manifold is restated here for convenience.
Definition 31.
Let $E$ be a locally free submodule of the $C^{\infty}$ module $\mathfrak{X}(M)$ of vector fields on $M$. By the Serre-Swan theorem, there is an $E$-tangent bundle ${}^{E}TM$ whose sections (locally) are sections of $E$, and an $E$-cotangent bundle ${}^{E}T^{*}M:=\left({}^{E}TM\right)^{*}$. We will call the global sections of $\Lambda^{p}\left({}^{E}T^{*}M\right)$ $E$-forms of degree $p$, and denote the space of all such sections by ${}^{E}\Omega^{p}(M)$. If $E$ satisfies the involutivity condition $[E,E]\subseteq E$, there is a differential $d:{}^{E}\Omega^{p}(M)\rightarrow{}^{E}\Omega^{p+1}(M)$ given by
${}^{E}d\eta\left(V_{0},\ldots,V_{p}\right)=$ $\sum_{i}(-1)^{i}V_{i}\left(\eta\left(V_{0},\ldots,\hat{V}_{i},\ldots,V_{p}\right)\right)$
$+\sum_{i<j}(-1)^{i+j}\eta\left(\left[V_{i},V_{j}\right],V_{0},\ldots,\hat{V}_{i},\ldots,\hat{V}_{j},\ldots,V_{p}\right).$
The cohomology of this complex is known as $E$-cohomology and is the same as Lie Algebroid Cohomology if we view this as a Lie Algebroid with anchor given by inclusion. If we have an $E$-closed nondegenerate $E$-$2$-form $\omega$, we call this an $E$-symplectic form and the triple $(E,M,\omega)$ an $E$-symplectic manifold. These are Symplectic Lie Algebroids. The cohomology of such an object in general is rather difficult to compute, however Miranda and Scott showed the following
Theorem 4.1 (Miranda and Scott [Miranda_2020]).
Let $M=\mathbb{R}^{2}$, and let $E$ be the involutive subbundle of $\mathfrak{X}(\mathbb{R}^{2})$ generated by
$v_{1}$ $=x\partial_{x}+y\partial_{y}$
$v_{2}$ $=-y\partial_{x}+x\partial y$
With dual bundle generated by
$v_{1}^{*}$ $=\frac{xdx+ydy}{x^{2}+y^{2}}$
$v_{2}^{*}$ $=\frac{-ydx+xdy}{x^{2}+y^{2}}$
Then the $E$-Cohomology or Lie Algebroid Cohomology is given by
${}^{E}H^{k}(M)=\begin{cases}\mathbb{R}\quad&i=0,2\\ \mathbb{R}^{2}&i=1\\ 0&i\geq 3\end{cases}$
Theorem 4.2.
The Symplectic Lie Algebroid defined by Miranda and Scott $(E\rightarrow\mathbb{R}^{2},\rho=\text{id},\omega=v_{1}^{*}\wedge v_{2}^{*})$ satisfies theorem 3.4
Proof.
We only need to check that the map $[L]:{}^{E}H^{0}(\mathbb{R}^{2})\rightarrow{}^{E}H^{2}(\mathbb{R}^{2})$ is an isomorphism, which amounts to showing that $\omega=v_{1}^{*}\wedge v_{2}^{*}$ is not an exact form. Suppose towards a contradiction that it is, i.e. there are $f,g\in C^{\infty}(\mathbb{R}^{2})$ such that $\omega={}^{E}d(fv_{1}^{*}+gv_{2}^{*})$. Unfolding the definition this implies that
$\omega=(v_{1}(g)-v_{2}(f))\omega$
Which means for all $(x,y)\in\mathbb{R}^{2}$ we have that
$v_{1}(g)-v_{2}(f)=1$
Or equivalently
$\left(\frac{\partial g}{\partial x}-\frac{\partial f}{\partial y}\right)x+\left(\frac{\partial g}{\partial x}+\frac{\partial f}{\partial x}\right)y=1$
However at $(0,0)$ this fails, so $\omega$ is not exact ∎
4.5. Kähler Lie Algebroids
There has been recent attention towards Kähler Lie Algebroids, see [hu2024kodairavanishingtheoremskahler]. By the Hard Lefschetz Theorem, Kälher manifolds are always Lefschetz manifolds. This relationship still holds true for a suitable class of Kähler Lie Algebroids
Definition 32.
A “Kähler Lie Algebroid” $(\mathcal{A}\rightarrow M,\rho,g)$ is a Lie Algebroid with a Hermitian metric $g$ whose associated canonical $(1,1)$-$\mathcal{A}$-form $\omega\in\Omega_{\mathcal{A}}^{1,1}(M)$ is closed
Corollary 4.2.1.
Kähler Lie Algebroids are Symplectic Lie Algebroids
Throughout this section, let $J$ be the associated complex structure for $\mathcal{A}$. Reconsidering the construction of $d_{\mathcal{A}}^{\star}$ with the presence of a complex structure reveals that
$Jd_{\mathcal{A}}^{\star}J^{-1}=d_{\mathcal{A}}^{\dagger}$
Where $d_{\mathcal{A}}^{\dagger}$ is the hermitian adjoint to $d_{\mathcal{A}}$, for details see [Shlomonotes].
Theorem 4.3.
Kähler Lie Algebroids which admit harmonic representatives for every $\mathcal{A}$-cohomology class in the Riemannian sense are Lefschetz Lie Algebroids
Proof.
It suffices to show that every cohomology class admits a symplectic harmonic representative. Fix any class $[\alpha]\in H_{\mathcal{A}}^{p,q}(M,\mathbb{C})$, where we’ve chosen $\alpha$ to be a harmonic representative. Then $d_{A}\alpha=d_{\mathcal{A}}^{\dagger}\alpha=0$. By our above observations we have
$0=d_{\mathcal{A}}^{\dagger}\alpha=Jd_{\mathcal{A}}^{\star}J^{-1}\alpha$
Implying that $(-1)^{p-q}d_{\mathcal{A}}^{\star}\alpha=0$, so $\alpha$ is symplectic harmonic ∎